### v2

# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

Abstract.
The purpose of this article is to establish a sharp version of Adams’ type inequality in a suitable higher order function space with singular weight in $\mathbb{R}^{n}$. In addition, we also provide the proof of a sharp singular concentration-compactness principle due to Lions’ as an improvement of this singular Adams’ inequality. It is well known that such types of inequalities can be classified in two different ways, for instance, critical sharp singular Adams’ type inequality (involving Sobolev full norm as constraint) and subcritical sharp singular Adams’ type inequality (involving Sobolev partial norm as constraint). Further, we shall establish that critical and subcritical sharp singular Adams’ type inequalities are surprisingly equivalent. To be more precisely, we also discuss about the asymptotic behavior of the lower and upper bounds of subcritical sharp singular Adams’ type inequality and establish a relation between the suprema of such types of critical and subcritical sharp inequalities. Despite this, we shall demonstrate a new compact embedding, which plays a crucial role in our arguments. Moreover, as an application of these results, by employing the mountain pass theorem, we study the existence of nontrivial solutions to a class of nonhomogeneous quasilinear elliptic equations involving $(p,\frac{n}{2})$-biharmonic operator with singular exponential growth as follows
$\Delta^{2}_{p}u+\Delta^{2}_{\frac{n}{2}}u=\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad\mathbb{R}^{n},$
with $n\geq 4$, $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and the nonlinear term $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function, which behaves like $\exp{(\alpha|s|^{\frac{n}{n-2}})}$ as $|s|\to~{}+\infty$ for some $\alpha>0$.
Key words and phrases:
$(p,\frac{n}{2})$-biharmonic operator, Adams’ inequality, singular exponential growth, mountain pass theorem
1991 Mathematics Subject Classification:
35A23,35B33,35A23,35J30,26D10
corresponding author
1. Introduction and main results
In this article, we investigate the existence of nontrivial solutions to a class of $(p,\frac{n}{2})$-biharmonic equations with singular exponential growth in $\mathbb{R}^{n}$. Specifically, we study the following equation
$\Delta^{2}_{p}u+\Delta^{2}_{\frac{n}{2}}u=\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad\mathbb{R}^{n},$ ($\boldsymbol{S}$)
with $n\geq 4$, $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $\Delta^{2}_{t}~{}\cdot:=\Delta(|\Delta\cdot|^{t-2}\Delta~{}\cdot)$ is a fourth order operator, which is known as the standard $t$-biharmonic operator for any $t\geq 1$. The nonlinearity $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function and has critical exponential growth at infinity, that is, it behaves like $\exp{(\alpha|s|^{\frac{n}{n-2}})}$ as $|s|\to~{}+\infty$ for some $\alpha>0$, which means that there exists a positive constant $\alpha_{0}$ such that
$\lim_{|s|\to+\infty}|g(x,s)|\operatorname{exp}(-\alpha|s|^{\frac{n}{n-2}})=\begin{cases}0&\text{if }\alpha>\alpha_{0},\\ +\infty&\text{if }\alpha<\alpha_{0},\end{cases}$
uniformly with respect to $x\in\mathbb{R}^{n}$. This notion of criticality is basically driven by the well known Moser-Trudinger type inequality. Throughout this article, without further mentioning, we assume that $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ satisfies the following properties.
  • ($g_{1}$)
    The map $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function, the map $\cdot\mapsto g(x,\cdot)$ is an odd function and $g(x,0)=0$ for all $x\in\mathbb{R}^{n}$. In addition, we also have $g(x,s)>0$ for all $(x,s)\in\mathbb{R}^{n}\times(0,+\infty)$ and $g(x,s)<0$ for all $(x,s)\in\mathbb{R}^{n}\times(-\infty,0)$.
  • ($g_{2}$)
    There exists constant $\tau>\max\{p^{\ast},\frac{n}{2}\}$ with $p^{\ast}:=\frac{np}{n-2p}$, so-called critical Sobolev exponent such that
    $\lim_{s\to 0}\frac{|g(x,s)|}{\quad|s|^{\tau-1}}=0\quad\text{uniformly with respect to}~{}~{}x\in\mathbb{R}^{n}.$
  • ($g_{3}$)
    There exists constant $\alpha_{0}>0$ such that
    $\lim_{|s|\to+\infty}\frac{|g(x,s)|}{\operatorname{exp}(\alpha|s|^{\frac{n}{n-2}})}=\begin{cases}0&\text{if }\alpha>\alpha_{0},\\ +\infty&\text{if }\alpha<\alpha_{0},\end{cases}\quad\text{uniformly with respect to}~{}~{}x\in\mathbb{R}^{n}.$
  • ($g_{4}$)
    There exists constant $\mu>\frac{n}{2}$ such that
    $0<\mu G(x,s):=\mu\int_{0}^{s}g(x,t)\,\mathrm{d}t\leq g(x,s)s\quad\text{ for all}\quad(x,s)\in\mathbb{R}^{n}\times(\mathbb{R}\setminus\{0\}).$
  • ($g_{5}$)
    There exists positive constants $s_{0}$ and $M_{0}$ such that for all $x\in\mathbb{R}^{n}$
    $0<G(x,s)\leq M_{0}|g(x,s)|\quad\text{for all}\quad|s|\geq s_{0}.$
  • ($g_{6}$)
    There exists constant $\vartheta>\max\{p^{\ast},\frac{n}{2}\}$ and $\lambda>0$ such that
    $G(x,s)\geq\lambda|s|^{\vartheta}\quad\text{ for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}.$
In addition, one can notice that, due to $(g_{2})$ and $(g_{3})$, for any fixed $q>\max\{p^{\ast},\frac{n}{2}\}$ and $\alpha>\alpha_{0}$, there exists $\zeta>0$ and a constant $D_{\zeta}>0$ depending upon $\zeta,~{}\alpha$ and $q$ such that
$|g(x,s)|\leq\zeta|s|^{\tau-1}+D_{\zeta}|s|^{q-1}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}$ (1.1)
and by using $(g_{4})$, we have
$|G(x,s)|\leq\zeta|s|^{\tau}+D_{\zeta}|s|^{q}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R},$ (1.2)
where $\Phi_{\alpha,j_{0}}:\mathbb{R}\to[0,+\infty)$ is a Young function, which is defined by
$\Phi_{\alpha,j_{0}}(s):=\operatorname{exp}(\alpha|s|^{\frac{n}{n-2}})-\sum_{j=0}^{j_{0}-1}\frac{\alpha^{j}|s|^{\frac{nj}{n-2}}}{j!}\quad\text{with}\quad j_{0}:=\left\lceil\frac{p^{\ast}(n-2)}{n}\right\rceil=\min\Big{\{}j\in\mathbb{N}:~{}j\geq\frac{p^{\ast}(n-2)}{n}\Big{\}},$ (1.3)
[…]
Example 1.1.
We remark that the function defined by
$G(x,s):=\lambda|s|^{\vartheta}\operatorname{exp}(\alpha_{0}|s|^{\frac{n}{n-2}})\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}$
with $n\geq 4$, $\lambda>0$, $0<\alpha_{0}<,\alpha$, $\vartheta>\tau,\mu>\max\{p^{\ast},\frac{n}{2}\}$, $g(x,s)=\frac{\partial G(x,s)}{\partial s}$ for all $(x,s)\in\mathbb{R}^{n}\times(\mathbb{R}\setminus\{0\})$, and $g(x,0)=0$ for all $x\in\mathbb{R}^{n}$, then $g$ satisfies all the hypotheses $(g_{1})-(g_{6})$.
To deal with nonlinear elliptic PDEs, it is worthwhile to mention that several authors work with Sobolev spaces of different orders. Moreover, to understand the basic features of the Sobolev spaces, one sees that for any $m\in\mathbb{N}$ and $p\geq 1$, the space $W^{m,p}(\mathbb{R}^{n})$ can be categorized in three different ways, notably
  • (a)
    the Gagliardo-Nirenberg-Sobolev case: $mp<n$,
  • (b)
    the Sobolev borderline case: $mp=n$, and
  • (c)
    the Morrey’s case: $mp>n$.
The Gagliardo-Nirenberg-Sobolev inequality asserts that the embedding $W^{m,p}(\mathbb{R}^{n})\hookrightarrow L^{r}(\mathbb{R}^{n})$ is continuous for all $r\in[p,\frac{np}{n-mp}]$ and $mp<n$. In this case, we can study the variational problems in the Sobolev spaces $W^{m,p}(\mathbb{R}^{n})$ with subcritical and critical polynomial growth, that is, the nonlinear term cannot exceed the polynomial of degree $\frac{np}{n-mp}$. Spite of this, the Sobolev borderline case is very special because $\frac{np}{n-mp}\to+\infty$ as $mp\to n$, which suggests us that $W^{m,p}(\mathbb{R}^{n})\subset L^{\infty}(\mathbb{R}^{n})$. Unfortunately, this only holds whenever $m=p=n=1$ and one knows it may fail for $n\geq 2$. In this instance, we note that every polynomial growth can be entertained, and thus, one may ask for a function with another kind of maximal growth. Moser-Trudinger and Adams’ inequalities play key roles in handling such difficulties. In this direction, for any bounded smooth domain $\Omega\subset\mathbb{R}^{n}~{}(n\geq 2)$ with $p=n$ and $m=1$, Trudinger [47] established that the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$ holds, where $L_{\Phi_{\alpha}}(\Omega)$ denotes the Orlicz space associated with the Young function $\Phi_{\alpha}(t)=\operatorname{\exp}(\alpha|t|^{\frac{n}{n-1}})-1$ for some constant $\alpha>0$.The sharpness of the exponent $\alpha$ was first studied by Moser [40] and proved a version of Moser-Trudinger inequality as follows
$\sup_{u\in W^{1,n}_{0}(\Omega),~{}\|\nabla u\|_{L^{n}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\operatorname{\exp}(\alpha|u|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(n)|\Omega|,~{}\forall~{}~{}\alpha\leq\alpha_{n},$ (1.4)
where $\alpha_{n}=n\omega_{n-1}^{\frac{1}{n-1}}$ with $\omega_{n-1}$ is the measure of the unit sphere in $\mathbb{R}^{n}$. In addition, the above supremum is no longer finite if $\alpha>\alpha_{n}$. Later, Adimurthi-Sandeep [3] generalized the result obtained by Moser in [40] and established a sharp singular Moser-Trudinger inequality with the same Dirichlet norm used in (1.4). Both nonsingular and singular Moser-Trudinger inequalities in the Euclidean space have been studied in [12, 43] in $\mathbb{R}^{2}$, then for higher dimensions in [20, 1, 37, 4, 18, 21], respectively. Such inequalities have several applications in conformal geometry and geometric analysis of PDEs; for a detailed study, we refer to [14, 9, 30].
Due to the loss of compactness in the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$, the main difficulty arises in studying variational problems is to prove Palais-Smale compactness condition. To avoid this, Lions’ [38] introduced a concentration-compactness principle, basically a generalization of the Moser-Trudinger inequality. It states that if $u_{k}\rightharpoonup u$ weakly in $W^{1,n}_{0}(\Omega)$ with $u\neq 0$, $\|\nabla u_{k}\|_{L^{n}(\Omega)}\leq 1$, $|\nabla u_{k}|^{n}\rightharpoonup\mu$ weakly in $\mathcal{M}(\overline{\Omega})$, where $\mathcal{M}(\overline{\Omega})$ denotes the space of Radon measure in $\mathbb{R}^{n}$, then
$\sup_{k\in\mathbb{N}}\int_{\Omega}\operatorname{\exp}(\alpha_{n}\wp|u_{k}|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(\wp,\Omega),~{}\forall~{}~{}\wp\in[1,\eta),$ (1.5)
where $\eta=(1-\|\nabla u^{\ast}\|^{n}_{L^{n}(\Omega)})^{-\frac{1}{n-1}}$ and $u^{\ast}$ is the spherical symmetric decreasing rearrangement of $u$. Later, in [53] the authors sharpened the constant Weierstrass-$\wp$ in (1.5) and established an improved version of Lions’ concentration-compactness principle in $W^{1,n}_{0}(\Omega)$. Similar results for $\Omega=\mathbb{R}^{n}$ can be found in [22, 50].
To the best of our knowledge, there has been significant progress in the Moser-Trudinger inequalities to analyse the existence and multiplicity of solutions to quasilinear elliptic equations involving $n$-Laplace operator as well as $(p,n)$-Laplace operator in whole space $\mathbb{R}^{n}$. For a detailed study, one may see [4, 49, 29, 13].
Let us introduce some notations to familiarise ourselves with higher-order Sobolev spaces. For any $m\in\mathbb{N}$ and $u\in C^{m}$, the class of $m$-th order differentiable function, we denote $\nabla^{m}$ as the $m$-th order gradient operator, which is given by
$\nabla^{m}u:=\begin{cases}\Delta^{\frac{m}{2}}u\quad\quad\quad\text{if}~{}m~{}\text{is even},\\ \nabla\Delta^{\frac{m-1}{2}}u~{}\quad\text{if}~{}m~{}\text{is odd}.\end{cases}$
Moreover, for $m<n$, the Sobolev space with homogenous Navier boundary condition on $\Omega\subset\mathbb{R}^{n}$ is denoted by $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ such that
$W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega):=\Big{\{}u\in W^{m,\frac{n}{m}}(\Omega):~{}\Delta^{j}u_{|\partial\Omega}=0~{}\text{in the sense of trace,}~{}0\leq j\leq\Big{[}\frac{m-1}{2}\Big{]}\Big{\}},$
where []delimited-[][$\cdot$][ ⋅ ] stands for greatest integer function (see Tarsi [46]). In addition, we can notice that $W^{m,\frac{n}{m}}_{0}(\Omega)\subsetneq W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ and the space $W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega)$ can be explicitly defined by
$W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega):=W_{0}^{1,\frac{n}{2}}(\Omega)\cap W^{2,\frac{n}{2}}(\Omega),\quad\forall~{}n>2.$
In literature, it was Adams [2], who first studied the Moser-Trudinger inequality that appears in (1.4) into higher order Sobolev spaces. He established that for $m\in\mathbb{N}$ with $m<n$ and $\Omega$ is any bounded domain in $\mathbb{R}^{n}$, then
$\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega)_{,}~{}\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})\,\mathrm{d}x\leq C(n,m)|\Omega|,~{}\forall~{}~{}\beta\leq\beta(n,m),$ (1.6)
where
$\beta(n,m):=\frac{n}{\omega_{n-1}}\begin{cases}\Big{[}\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big{(}\frac{m+1}{2}\big{)}}{\Gamma\big{(}\frac{n-m+1}{2}\big{)}}\Big{]}^{\frac{n}{n-m}}\quad\text{if}~{}m~{}\text{is odd},\\ \Big{[}\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big{(}\frac{m}{2}\big{)}}{\Gamma\big{(}\frac{n-m}{2}\big{)}}\Big{]}^{\frac{n}{n-m}}\quad\quad\text{if}~{}m~{}\text{is even}.\end{cases}$
Despite this, the constant $\beta(n,m)$ is optimal in the sense that the supremum (1.6) will become infinite if $\beta>\beta(n,m)$. Moreover, (thanks to [46]) the Adams’ inequality also valid whenever $W^{m,\frac{n}{m}}(\Omega)$ is replaced by the larger space $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$. Subsequently, we remark here that the above Adams’ inequality (1.6) has been extended in many directions. For example, in the bounded domain, the following inequality with singular weight has been proved in [32, Lam-Lu],
$\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega),~{}\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\frac{\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty$ (1.7)
for all $0\leq\beta\leq\beta_{\alpha,n,m}=(1-\frac{\alpha}{n})\beta(n,m)$ and $\alpha\in[0,n)$. In addition, $\beta_{\alpha,n,m}$ is sharp, that is, the supremum in (1.7) is infinite if $\beta>\beta_{\alpha,n,m}$. It was also noted that if $m$ is an even integer, then this inequality is also true for the Sobolev space $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$. Next, for any $u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n})$, we define
$\|u\|_{m,\frac{n}{m}}:=\begin{cases}\|(I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}m~{}\text{is even},\\ \Big{(}\|(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}\Big{)}^{\frac{m}{n}}\quad\quad\quad\text{if}~{}m~{}\text{is odd},\end{cases}$
where $\|\cdot\|_{p}$ denotes $L^{p}(\mathbb{R}^{n})$ norm for $p\in[1,+\infty)$. Then the related inequalities corresponding to (1.6) when $\Omega=\mathbb{R}^{n}$ have been first proposed by [44, Ruf-Sani] for any positive even integer $m$. After that, it was again generalized in [31, Lam-Lu] for any arbitrary integer $m\geq 1$. Their combined results can be stated as follows
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|u\|_{m,\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\Phi(\beta(n,m)|u|^{\frac{n}{n-m}})\,\mathrm{d}x<+\infty,$ (1.8)
where
$\Phi(t):=\operatorname{exp}(t)-\sum_{j=0}^{j_{\frac{n}{m}}-2}\frac{t^{j}}{j!}\quad\text{with}\quad j_{\frac{n}{m}}:=\min\Big{\{}j\in\mathbb{N}:~{}j\geq\frac{n}{m}\Big{\}}.$
In addition, the constant $\beta(n,m)$ in the above supremum is sharp; that is, the supremum in (1.8) is infinity if we replace $\beta(n,m)$ by any $\beta>\beta(n,m)$. Further improvement of Adams’ inequalities (1.8) have been studied in [52] and proved that (1.8) still holds under the constraint:
$\big{\{}u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}):~{}\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}\leq 1~{}\text{for all}~{}\tau>0\big{\}}.$
However, they did not sharpen it, and we believe it can be sharpened by utilizing a similar strategy that appears in [41, Theorem 1.1].
It was Yang [48] who first extended the results in [44] and proved a version of subcritical singular Adams’ inequality in dimension four by introducing the Sobolev full norm. Basically, he established that
$\sup_{u\in W^{2,2}(\mathbb{R}^{4}),~{}\int_{\mathbb{R}^{4}}(|\Delta u|^{2}+\tau|\nabla u|^{2}+\sigma|u|^{2})~{}\mathrm{d}x~{}\leq 1}~{}\int_{\mathbb{R}^{4}}\frac{\operatorname{\exp}(\alpha u^{2})-1}{|x|^{\beta}}\,\mathrm{d}x<+\infty,~{}\forall~{}\alpha\in\Big{(}0,32\pi^{2}\Big{(}1-\frac{\beta}{4}\Big{)}\Big{)}$ (1.9)
for all constants $\tau,\sigma>0$ and $\beta\in[0,4)$. Further, when $\alpha>32\pi^{2}\big{(}1-\frac{\beta}{4}\big{)}$, the supremum (1.9) is no longer finite. Note that it has remained an open question to the readers whether the above inequality (1.9) still holds for the critical case, that is, $\alpha=32\pi^{2}\big{(}1-\frac{\beta}{4}\big{)}$. Then, again the singular Adams’ inequality (1.7) was fully extended into $\mathbb{R}^{n}$ if the Dirchlet norm $\|\nabla^{m}\cdot\|_{L^{\frac{n}{m}}(\Omega)}$ is replaced by $\|(\tau I-\Delta)^{\frac{m}{2}}\cdot\|_{\frac{n}{m}}$ for any $\tau>0$ (see [33]). More precisely, they proved that
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|(\tau I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~{}\forall~{}\beta\leq\beta_{\alpha,n,m},$ (1.10)
for all $m\in\mathbb{N}$ with $m<n$ and $\alpha\in[0,n)$. Consequently, if $\beta>\beta_{\alpha,n,m}$, then the supremum (1.10) is no longer finite. Note that in the second order Sobolev space $W^{2,\frac{n}{2}}(\mathbb{R}^{n})$ with $n\geq 4$, the sharp singular Adams’ inequality was investigated in [33] under the following constraint
$\big{\{}u\in W^{2,\frac{n}{2}}(\mathbb{R}^{n}):~{}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}+\tau\|u\|^{\frac{n}{2}}_{\frac{n}{2}}\leq 1~{}\text{for all}~{}\tau>0\big{\}}.$
Later, this result was again generalized in [52, Theorem 1.1] into the arbitrary Sobolev space $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$. Explicitly, they proved a strengthened version of the singular Adams’ inequality as follows: let $\tau>0$, then
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~{}\forall~{}\beta\leq\beta_{\alpha,n,m},$ (1.11)
for all $m\in\mathbb{N}$ with $m<n$ and $\alpha\in[0,n)$. It has been observed that the inequality (1.11) was not sharpened yet. Moreover, the existence and nonexistence of extremal functions for sharp Adams’ inequalities can be found in some recent articles (see [39, 17, 16, 51]).
In recent years, Adams’ type inequalities have been widely studied by many authors across diverse domains such as Hyperbolic spaces, Lorentz spaces, CR spheres, compact Riemannian manifolds, and so on. In this context, we cite some delightful works captured in [28, 27, 26]. For the applications of Adams’ inequalities in the analysis of elliptic PDEs, we also refer to some notable articles [24, 15, 8].
It is worth mentioning that the concentration-compactness principle was first studied in higher-order Sobolev spaces in [23, Theorem 1]. Later, on exploiting the Hilbert structure of $W^{2,2}(\mathbb{R}^{4})$, in [15] authors established a sharp singular concentration-compactness principle and as an application they studied a class of biharmonic equations in dimension four. Then, it was fully extended into whole space $\mathbb{R}^{n}$ in [41]. His result can be read as follows: let $u_{k}\rightharpoonup u$ weakly in $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$ with $u\neq 0$, $\|u_{k}\|_{m,\frac{n}{m}}\leq 1$ and
$1<\wp<Q_{n,m}(u):=\Big{(}1-\|u\|^{\frac{n}{m}}_{m,\frac{n}{m}}\Big{)}^{-\frac{m}{n-m}}\quad(~{}Q_{n,m}(u)=+\infty\quad\text{when}\quad\|u\|_{m,\frac{n}{m}}=1),$
then it holds that
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\Phi(\wp\beta(n,m)|u_{k}|^{\frac{n}{n-m}})\,\mathrm{d}x<+\infty.$
In addition, he also proved that the above result holds whenever $\|\cdot\|_{m,\frac{n}{m}}$ is replaced by the Ruf norm on $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$, which is defined by
$\|u\|_{\text{Ruf}}:=\big{(}\|u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}\big{)}^{\frac{m}{n}},~{}\forall~{}u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}).$
Moreover, we also mention that the asymptotic behavior of the supremum for the subcritical sharp Moser-Trudinger inequalities in $\mathbb{R}^{n}$ with different function spaces can be found in the recent works [36, 35, 45]. In addition, it was also proved in these articles that the critical and subcritical sharp Moser-Trudinger inequalities are indeed equivalent. Similar results for higher order Sobolev spaces in $\mathbb{R}^{n}$ can be seen in [34] and references therein.
To study ($\boldsymbol{S}$), we first define some suitable function spaces. Notice that due to the appearance of two operators in ($\boldsymbol{S}$), we work with the function space
$E:=\Big{\{}u\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})\colon~{}\int_{\mathbb{R}^{n}}|\Delta u|^{p}\,\mathrm{d}x<+\infty,~{}\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}}\,\mathrm{d}x<+\infty\Big{\}},$
where $1<p<\frac{n}{2}$ with $n\geq 4$, which is the completion of the $C_{0}^{\infty}(\mathbb{R}^{n})$ with respect to the following norm
$\|u\|:=\Big{(}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}+\|\Delta u\|^{\frac{n}{2}}_{p}\Big{)}^{\frac{2}{n}},~{}\forall~{}u\in E,$
[…]
$\|u\|\leq\|u\|_{E}\leq 2^{\frac{n-2}{n}}\|u\|,~{}\forall~{}u\in E.$ (1.12)
Using standard arguments, we remark that $(E,\|\cdot\|_{E})$ is a uniform convex, reflexive and separable Banach space. Moreover, since reflexivity is preserved under equivalent norms, we infer that $(E,\|\cdot\|)$ is also a reflexive Banach space. In addition, the space $D^{2,p}(\mathbb{R}^{n})$ is defined as the completion of $C_{0}^{\infty}(\mathbb{R}^{n})$ equipped with the norm
$\|u\|_{D^{2,p}}:=\|\Delta u\|_{p},~{}\forall~{}u\in D^{2,p}(\mathbb{R}^{n}).$
Due to the Sobolev embedding, one can see that the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ hold.
For any $\gamma\in[0,n)$, we define the space $L^{r}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$, consisting of all real-valued measurable functions, with $|u|^{r}|x|^{-\gamma}\in L^{1}(\mathbb{R}^{n})$, equipped with the norm
$\|u\|_{r,\gamma}:=\bigg{(}\int_{\mathbb{R}^{n}}|u|^{r}|x|^{-\gamma}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}.$
Moreover, throughout this article, we have the following notations:
[…]
  • $o_{k}(1)$ denotes the real sequence converges to zero as $k\to\infty$,
[…]
  • $B_{r}(y)$ denotes an open ball centered at $y\in\mathbb{R}^{n}$ with radius $r>0$ and $B_{r}=B_{r}(0)$,
  • $f\lesssim~{}(\gtrsim)~{}g$, which means $f\leq(\geq)~{}Cg$ for some suitable constant $C>0$. We shall also write $f\sim g$ to denote that $f\lesssim g$ and $f\gtrsim g$,
  • $f\underset{\approx}{<}g$ means $f$ is close enough to $g$, $f\thickapprox g$ stands for $f$ is approximately equal to function $g$ and $O(f)$ means some constant multiple of $f$,
  • $|A|$ represents $n$-dimensional Lebesgue measure for $A\subset\mathbb{R}^{n}$; and its complement by $A^{c}$.
We would like to emphasize the important aspects and novelties of this article below:
  • (a)
    In higher order Sobolev spaces, the Pólya-Szegő types inequalities don’t hold, which causes us to face several barriers in proving the singular Adams’ type inequality in $E$ and the associated concentration-compactness principle, as can be shown in Theorem 1.3 and Theorem 1.4.
  • (b)
    Due to the critical exponential growth of the nonlinearity, there is a healthy competition between the nonhomogeneous $(p,\frac{n}{2})$-biharmonic operator and the nonlinear term because the embedding $E\hookrightarrow L_{\Phi_{\alpha,j_{0}}}(\mathbb{R}^{n})$ is not compact, where $\Phi_{\alpha,j_{0}}(\cdot)$ is the Young function defined in (1.3). Therefore, proving the compactness of Palais-Smale sequences for the variational energy associated with our main problem becomes challenging.
  • (c)
    The study of existence of solutions to our main problem relies on the compact embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$, thanks to Theorem 1.7. It follows that our method is different from $\gamma=0$.
  • (d)
    Owing to the fact that $E$ is not a Hilbert space, for any bounded Palais-Smale sequence $\{u_{k}\}_{k}\subset E$ with $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$, we cannot directly get
    $|\Delta u_{k}|^{t-2}\Delta u_{k}\rightharpoonup|\Delta u|^{t-2}\Delta u\quad\text{in}\quad L^{\frac{t}{t-1}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty$
    for $t\in\{p,\frac{n}{2}\}$. It indicates that more delicate analysis is required to prove
    $\Delta u_{k}\to\Delta u\quad\text{a.e. in}\quad\mathbb{R}^{n}\quad\text{as}\quad k\to\infty.$
  • (e)
    It is well known that for any $u\in E$, there is no guarantee that $|u|,~{}u^{\pm}\in E$, where $u^{\pm}=\max\{\pm u,0\}$ and thus we shall unable to get positive solutions to our main problem.
  • (f)
    Sophisticated methods, such as variational and topological tools, are used in the proofs.
Motivated by the above-mentioned works and primarily inspired by the works in [48, 33, 41, 15, 13, 5, 34], we establish a new version of sharp singular Adams’ inequality in the space $E$, its concentration-compactness principle, and as a byproduct of these inequalities, we study a class of elliptic pde involving $(p,\frac{n}{2})$-biharmonic operator in the entire space $\mathbb{R}^{n}$. The following defines the weak formulation of our main problem.
Definition 1.2.
We say that $u\in E$ is a weak solution to the problem ($\boldsymbol{S}$), if there holds
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u~{}\Delta v~{}\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u~{}\Delta v~{}\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}v\,\mathrm{d}x,~{}\forall~{}v\in E.$
Briefly speaking, as our interest is to study the problem ($\boldsymbol{S}$) in the space $E$, it should pointed out that, unfortunately, we cannot use the singular Adams’ inequality that appeared in [33, Theorem 1.5]. Hence, we need to demonstrate a version of such inequality.
Theorem 1.3 (Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Then for all $0\leq\alpha\leq\beta_{\gamma,n}:=\big{(}1-\frac{\gamma}{n}\big{)}\beta(n,2)$ and $u\in E$, there holds
$\sup_{u\in E,~{}\|u\|\leq 1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty,$ (1.13)
where $\Phi_{\alpha,j_{0}}(\cdot)$ is defined in (1.3) and $\beta(n,2)=n[(n-2)\omega^{\frac{2}{n}}_{n-1}]^{\frac{n}{n-2}}$ with $\omega_{n-1}$ stands for the measure of the unit sphere in $\mathbb{R}^{n}$. Further, the constant $\beta_{\gamma,n}$ is sharp, that is, if $\alpha>\beta_{\gamma,n}$, then the supremum (1.13) is infinite.
As an improvement of the above singular Adams’ inequality, basically inspired by the notable works of do Ó-Macedo [23] and Nguyen [41], we shall prove a sharp singular concentration-compactness principle as follows.
Theorem 1.4 (Concentration-Compactness Principle).
Suppose that $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Let $\{u_{k}\}_{k}\subset E$ be a sequence such that $\|u_{k}\|=1$ and $u_{k}\rightharpoonup u\neq 0$ in $E$ as $k\to\infty$, then for all $\ell$ satisfying
$0<\ell<L_{n}(u):=\big{(}1-\|u\|^{\frac{n}{2}}\big{)}^{-\frac{2}{n-2}}\quad\quad(~{}L_{n}(u)=+\infty\quad\text{whenever}\quad\|u\|=1),$
we have
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$ (1.14)
Finally, the constant $L_{n}(u)$ is sharp in the sense that the supremum (1.14) is infinite for $\ell>L_{n}(u)$.
The following natural question is still open at this time:
[…]
The next result says about the lower and upper bounds of the subcritical sharp singular Adams’ type inequality in the space $E$ asymptotically. In fact, we illustrate the subsequent theorem.
Theorem 1.5 ( Subcritical Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$, $n\geq 4$ and $0\leq\ell<\beta(n,2)$. Then there holds
$\text{ATSC}(\ell,\gamma):=\sup_{u\in E,~{}\|\Delta u\|_{\frac{n}{2}}\leq 1}\cfrac{1}{\|\Delta u\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
The constant $\beta(n,2)$ is sharp in the sense that $\text{ATSC}(\beta(n,2),\gamma)=+\infty$. Also, there exists two positive constants $c(n,\gamma,p)$ and $C(n,\gamma,p)$ such that whenever $l\underset{\approx}{<}\beta(n,2)$,
$\cfrac{c(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}\leq\text{ATSC}(\ell,\gamma)\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}.$
In the following theorem, we first describe a necessary and sufficient condition for the boundedness of the critical sharp singular Adams’ inequality in the space $E$. Finally, we shall demonstrate the equivalence of critical and subcritical sharp singular Adams’ inequalities.
Theorem 1.6 ( Critical Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$, $n\geq 4$ and $a,b>0$. Then there holds
$\text{ATC}_{a,b}(\gamma):=\sup_{u\in E,~{}\|\Delta u\|^{a}_{\frac{n}{2}}+\|\Delta u\|^{b}_{p}\leq 1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty\iff b\leq\frac{n}{2}.$
The constant $\beta(n,2)$ is sharp in the sense that the above supremum is infinite when $\beta(n,2)$ is replaced by a larger constant. In addition, the following identity holds
$\text{ATC}_{a,b}(\gamma)=\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$ (1.15)
Moreover, there hold $\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma)<+\infty$ and
$\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma)=\sup_{\ell\in(0,\beta(n,2))}\bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}}{\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}}\Bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$
To state our main result, we need the following compact embedding, which play crucial rule in the study of the problem ($\boldsymbol{S}$). Its statement can be read in the following way.
Theorem 1.7.
Assume that $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $n\geq 4$. Then the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ is compact for all $\varrho\geq p^{\ast}$, where $p^{\ast}:=\frac{np}{n-2p}$, so-called critical Sobolev exponent.
Now, we can state the main result of this article as follows.
Theorem 1.8.
Suppose that $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $n\geq 4$. Let the hypotheses $(g_{1})-(g_{6})$ be satisfied. Moreover, let there exists $\lambda_{0}$ be sufficiently large enough such that $(g_{6})$ holds for $\lambda\geq\lambda_{0}$, then the problem ($\boldsymbol{S}$) has a nontrivial weak solution.
The structure of this article is organized in the following manner: Section 2 is dedicated to some preliminary results, which are needed in our proofs. In Section 3, we shall prove Theorem 1.3 and Theorem 1.4, respectively. Section 4 is devoted to the proof of Theorem 1.5 and Theorem 1.6, whereas in Section 5, we shall provide the proof of Theorem 1.7. Finally, due to the applications of these results, and using the mountain pass theorem, in Section 6, we prove the main result of the article, that is, Theorem 1.8.
2. Preliminary results
In this section, we shall discuss the basic properties of Schwarz symmetrization and some useful lemmas. In this regards, we refer to [7, 10]. First, we introduce about the symmetrization of Lebesgue measurable set $S$ in $\mathbb{R}^{n}$. Let $S^{\ast}$ be the open ball centered at the origin of $\mathbb{R}^{n}$ such that $|S^{\ast}|=|S|$. Further, we denote $|S|=0$ implies $S^{\ast}$ is the empty set, $|S|=+\infty$ implies $S^{\ast}=\mathbb{R}^{n}$ and $0<|S|<+\infty$ implies $S^{\ast}=B_{R}(0)$ such that $|S^{\ast}|=|S|=\sigma_{n}R^{n}$, where $\sigma_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $R=R(S)$ is the volume radius of $S$. It is well known that $S^{\ast}$ is called the symmetric decreasing rearrangement of $S$. Let $f\colon S\to\mathbb{R}$ be a measurable function vanishing at infinity, which means that
|$|\{x\in S\colon~{}|f(x)|>t\}|=\int_{\{x\in S\colon~{}|f(x)|>t\}}\mathrm{d}x<+\infty,~{}\forall~{}t>0.$
Its distribution function $\mu_{f}\colon[0,+\infty)\to[0,+\infty]$ is given by $\mu_{f}(t)=|\{x\in S\colon~{}|f(x)|>t\}|$ for all $t\geq 0$. Note that $\mu_{f}$ is always nonnegative, decreasing and right continuous function on $[0,+\infty)$. Moreover, the unidimensional rearrangement of $f$ is a function $f^{\sharp}\colon[0,|S|]\to[0,+\infty]$, which is defined by
$f^{\sharp}(0):=\text{ess sup}f\quad\text{and}\quad f^{\sharp}(s):=\inf\{t\geq 0\colon~{}\mu_{f}(t)<s\},~{}\forall~{}s\in(0,|S|].$
In fact, it has been observed by Bennett-Sharpley [10] that the following identity also holds
$f^{\sharp}(s)=m_{\mu_{f}}(s)=\sup\{t\geq 0:~{}\mu_{f}(t)>s\},~{}\forall~{}s\geq 0,$
where $m_{\mu_{f}}$ is the distribution function of $\mu_{f}$. Besides this, $f^{\sharp}$ is a nonnegative, decreasing and right continuous function on $[0,+\infty)$. The maximal function of $f^{\sharp}$ is denoted by the symbol $f^{\sharp\sharp}$, and given by
$f^{\sharp\sharp}(s):=\frac{1}{s}\int_{0}^{s}f^{\sharp}(t)\,\mathrm{d}t,~{}\forall~{}~{}s>0.$ (2.1)
In addition, $f^{\sharp\sharp}$ is nonnegative, decreasing and continuous on $(0,+\infty)$ and $f^{\sharp}\leq f^{\sharp\sharp}$. Now we denote the function $f^{\ast}\colon S^{\ast}\to[0,+\infty]$ as the Schwarz symmetrization of $f$, alternatively known as the spherically symmetric decreasing rearrangement of $f$, and defined by
$f^{\ast}(x):=f^{\sharp}(\sigma_{n}|x|^{n}),~{}\forall~{}x\in S^{\ast}.$
In particular, $f^{\ast}$ is a radially symmetric nonnegative nonincreasing function. Moreover, if $\Psi\colon\mathbb{R}\to\mathbb{R}$ is nondecreasing, then we have
$(\Psi(f))^{\ast}=\Psi(f^{\ast}).$ (2.2)
It has been observed that $f^{\ast}$ and $|f|$ are equimeasurable, that is, $\{x\colon~{}f^{\ast}>t\}=\{x\colon~{}|f|>t\}^{\ast}$ and $(\chi_{A})^{\ast}=\chi_{A^{\ast}}$, where $\chi_{A}$ is the indicator function associate with measurable set $A\subset\mathbb{R}^{n}$. Moreover, if $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is a measurable function vanishing at infinity, we can deduce from the Layer cake representation that
$f^{\ast}(x):=\int_{0}^{\infty}\chi_{\{x\colon~{}|f|>t\}^{\ast}}(x)\,\mathrm{d}t,~{}\forall~{}x\in\mathbb{R}^{n}.$
In addition, we also have $\|f\|_{p}=\|f^{\ast}\|_{p}$, $\forall~{}p\in[1,+\infty)$ and the following Hardy-Littlewood inequality holds.
Lemma 2.1.
[7, Corollary 2.16] Suppose that $f$ and $g$ are two nonnegative measurable functions on $\mathbb{R}^{n}$ with $\mu_{f}(t)<+\infty$ and $\mu_{g}(t)<+\infty$ for all $t>0$. Then
$\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x\leq\int_{\mathbb{R}^{n}}f^{\ast}(x)g^{\ast}(x)\,\mathrm{d}x$ (2.3)
with the understanding that when the left side is infinite, so is the right side. Moreover, if $\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x<+\infty$, then the equality in (2.3) holds if and only if |$|\{(x,y)\in\mathbb{R}^{2n}\colon~{}f(x)<(f(y)~{}\text{and}~{}g(x)>g(y)\}|=0.$
We highlight that the following lemmas are essential for studying the subsequent sections.
Lemma 2.2.
[23, Lemma 2] Let $S\subset\mathbb{R}^{n}$ be an open set and $f_{k},f\in L^{\wp}(S)$ such that $f_{k}\rightharpoonup f$ in $L^{\wp}(S)$ as $k\to\infty$ for any $\wp>1$. Then up to a subsequence $f_{k}^{\sharp}:=h_{k}\to h$ as $k\to\infty$ a.e. in $[0,|S|]$ for some $h\in L^{\wp}([0,|S|])$ with $\|h\|_{L^{\wp}([0,|S|])}\geq\|f^{\sharp}\|_{L^{\wp}([0,|S|])}$.
The following result is a crucial tool for proving Theorem 1.4.
Lemma 2.3.
[41, Proposition 2.2] Let $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ and $-\Delta u=f\in C^{\infty}_{0}(\mathbb{R}^{n})$, then there exists a constant $D(n)$ depends upon only $n$ such that
$u^{\sharp}(s_{1})-u^{\sharp}(s_{2})\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{f^{\sharp\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+D(n)\|f\|_{\frac{n}{2}},~{}\forall~{}~{}0<s_{1}<s_{2}<+\infty.$
Lemma 2.4.
[11, Lemma A.VI] If $u\in L^{\wp}(\mathbb{R}^{n}),~{}\wp\geq 1$, is a radial nonincreasing function, then there holds
$|u(x)|\leq|x|^{-\frac{n}{\wp}}\Big{(}\frac{n}{\omega_{n-1}}\Big{)}^{\frac{1}{\wp}}\|u\|_{\wp},~{}\forall~{}x\in\mathbb{R}^{n}\setminus\{0\}.$
Lemma 2.5.
[5, Proposition 1] Let $1<p<\frac{n}{2}$ with $n\geq 4$, then the embedding $E\hookrightarrow L^{\theta}(\mathbb{R}^{n})$ is continuous for all $\theta\in[p^{\ast},+\infty)$. In addition, the embedding $E\hookrightarrow L^{\theta}_{\textit{loc}}(\mathbb{R}^{n})$ is compact for all $\theta\in[1,+\infty)$.
The following corollary follows from [49, Lemma 2.1].
Corollary 2.6.
The function $s\mapsto\operatorname{\exp}(s)-\sum_{j=0}^{j_{0}-1}\frac{s^{j}}{j!}$ is continuous, increasing and convex in $[0,+\infty)$, where $j_{0}$ is defined in (1.3). Moreover, for any $\alpha>0$, $\wp>1$ and $n\geq 4$, we have
$\bigg{(}\operatorname{exp}(\alpha|s|^{\frac{n}{n-2}})-\sum_{j=0}^{j_{0}-1}\frac{\alpha^{j}}{j!}|s|^{\frac{nj}{n-2}}\bigg{)}^{\wp}\leq\operatorname{exp}(\alpha\wp|s|^{\frac{n}{n-2}})-\sum_{j=0}^{j_{0}-1}\frac{\alpha^{j}\wp^{j}}{j!}|s|^{\frac{nj}{n-2}},~{}\forall~{}s\in\mathbb{R}.$
Lemma 2.7.
[25, Theorem 1.2 and Lemma 3.5] If $q\in[1,+\infty]$, $r\in[1,+\infty)$, $j,~{}k\in\mathbb{N}$, $j<k$ such that $\frac{1}{p}=\frac{j}{kr}+\frac{k-j}{kq}$, then there exists a constant $C>0$ independent of $u$ such that
$\|\nabla^{j}u\|_{p}\leq C\|\nabla^{k}u\|^{\frac{j}{k}}_{r}\|u\|^{1-\frac{j}{k}}_{q},~{}\forall~{}u\in L^{q}(\mathbb{R}^{n})\cap W^{k,r}(\mathbb{R}^{n}).$
In particular, for $1<p<\frac{n}{2}$ with $n\geq 4$ and $t\in\{p,\frac{n}{2}\}$, there holds
$\|\nabla u\|_{t}\leq C\|\Delta u\|^{\frac{1}{2}}_{t}\|u\|^{\frac{1}{2}}_{t},~{}\forall~{}u\in W^{2,t}(\mathbb{R}^{n}).$
3. Proof of Theorem 1.3 and Theorem 1.4
In this section, our aim is to prove Theorem 1.3 and Theorem 1.4. To establish Theorem 1.3, we shall use a rearrangement-free approach introduced by Lam-Lu [33], whereas the proof of Theorem 1.4 relies on conventional Schwarz symmetrization techniques.
Proof of Theorem 1.3..
Invoking the fact that $\Phi_{\alpha,j_{0}}(s)=\Phi_{\alpha,j_{0}}(|s|)$ and exploiting the denseness of $C^{\infty}_{0}(\mathbb{R}^{n})$ in $E$, without loss of generality, we can take that $u\in C^{\infty}_{0}(\mathbb{R}^{n})\setminus\{0\}$ with $u\geq 0$ and $\|u\|\leq 1$. Define the following
$H(u):=2^{-\frac{4}{n(n-2)}}\|\Delta u\|_{p}\quad\text{and}\quad~{}S:=\{x\in\mathbb{R}^{n}\colon~{}u(x)>H(u)\}.$
It follows at once that
$H(u)<1\quad\text{and}\quad|S|\leq 2^{\frac{4p^{\ast}}{n(n-2)}}C,$ (3.1)
where $C>0$ is a suitable constant depending only on $n$ and $p$, thanks to the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$. This shows that $S$ is a bounded domain. In addition, we can notice that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x=\int_{S}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\mathbb{R}^{n}\setminus S}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x:=J_{1}+J_{2}.$
Initially, we shall evaluate $J_{2}$. By using (3.1), we have $\mathbb{R}^{n}\setminus S\subset\{x\in\mathbb{R}^{n}:~{}u(x)<1\}$. Due to Lemma 2.5 and the fact that $\|u\|\leq 1$, we obtain
$J_{2}$$\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1,~{}|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1,~{}|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\leq C(\alpha,\gamma,n,p),$ (3.2)
where $C(\alpha,\gamma,n,p)>0$ is a constant depending only on $\alpha$, $\gamma$, $n$ and $p$. In order to evaluate $J_{1}$, we first denote $w(x):=u(x)-H(u)$ in $S$. It is easy to see that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$. As a result, we have the following estimates
$|u|^{\frac{n}{n-2}}$ $\leq(|w|+H(u))^{\frac{n}{n-2}}\leq|w|^{\frac{n}{n-2}}+\frac{n}{n-2}2^{\frac{n}{n-2}-1}\big{[}|w|^{\frac{n}{n-2}-1}H(u)+|H(u)|^{\frac{n}{n-2}}\big{]}$
$\leq|w|^{\frac{n}{n-2}}+\frac{n}{n-2}2^{\frac{n}{n-2}-1}\Bigg{[}\frac{|w|^{\frac{n}{n-2}}|H(u)|^{\frac{n}{2}}}{\frac{n}{2}}+\frac{n-2}{n}+|H(u)|^{\frac{n}{n-2}}\Bigg{]}$
$\leq\bigg{[}1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg{]}|w|^{\frac{n}{n-2}}+D_{n},$
we thank to the well known Young’s inequality and the following elementary inequality of calculus
$(a+b)^{\wp}\leq a^{\wp}+\wp 2^{\wp-1}(a^{\wp-1}b+b^{\wp}),~{}\forall~{}\wp\geq 1;~{}a,b\geq 0.$
Set
$\xi(x):=\bigg{[}1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg{]}^{\frac{n-2}{n}}w(x)\quad\text{in}~{}~{}S.$
Observe that $\xi\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and thus $|u|^{\frac{n}{n-2}}\leq|\xi|^{\frac{n}{n-2}}+D_{n}$ in $S$. In addition, by direct calculation, we get
$\|\Delta\xi\|^{\frac{n}{2}}_{L^{\frac{n}{2}}(S)}\leq\bigg{[}1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg{]}^{\frac{n-2}{2}}\big{(}1-\|\Delta u\|^{\frac{n}{2}}_{p}\big{)}.$
The above inequality together with the fact that $(1-x)^{\wp}\leq 1-\wp x$ for all $x\in[0,1]$ and $\wp\in(0,1]$ imply that
$\|\Delta\xi\|^{\frac{n}{n-2}}_{L^{\frac{n}{2}}(S)}$ $\leq\bigg{[}1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg{]}\Big{(}1-{\frac{2}{n-2}}\|\Delta u\|^{\frac{n}{2}}_{p}\Big{)}$
$\leq\Big{(}1+{\frac{2}{n-2}}\|\Delta u\|^{\frac{n}{2}}_{p}\Big{)}\Big{(}1-{\frac{2}{n-2}}\|\Delta u\|^{\frac{n}{2}}_{p}\Big{)}$
$=1-\frac{4}{(n-2)^{2}}\|\Delta u\|^{n}_{p}\leq 1.$
It follows that $\xi\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and $\|\Delta\xi\|_{L^{\frac{n}{2}}(S)}\leq 1$. Consequently, by employing the singular Adams’ inequality established in [32, Theorem 1.2] with homogeneous Navier boundary condition, we obtain
$J_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S}\frac{\operatorname{\exp}(\beta_{\gamma,n}(|\xi|^{\frac{n}{n-2}}+D_{n}))}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\beta_{\gamma,n}D_{n})C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq C(\gamma,n,p),$ (3.3)
where $C(\gamma,n,p)>0$ is a constant depending only on $\gamma$, $n$ and $p$. In virtue of (3) and (3), we infer that (1.13) holds.
Next, to check the validity of the sharpness of $\beta_{\gamma,n}$, inspired by Lu-Yang [39], we can define the following sequence
$\xi_{k}(x):=\begin{cases}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}|x|^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\quad\text{if}~{}|x|\in[0,k^{-\frac{1}{n}}],\\ n\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\ln\big{(}\frac{1}{|x|}\big{)}\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in[k^{-\frac{1}{n}},1],\\ \eta_{k}(x)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in[1,+\infty),\end{cases}$ (3.4)
where $\{\eta_{k}\}_{k}$ is a sequence of radially smooth functions such that
$\text{supp}(\eta_{k})\subset\{x\in\mathbb{R}^{n}\colon~{}1<|x|<2\}\quad\text{and}\quad\frac{\partial\eta_{k}}{\partial\nu}\Big{|}_{\partial B_{1}}=n\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}.$
Moreover, $\eta_{k}$ and $\Delta\eta_{k}$ are all of $O((\ln k)^{-\frac{2}{n}})$. By direct calculation, one can notice that
$\Delta\xi_{k}(x):=\begin{cases}-n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in(0,k^{-\frac{1}{n}}),\\ -n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}|x|^{-2}\quad\quad\text{if}~{}|x|\in(k^{-\frac{1}{n}},1),\\ \Delta\eta_{k}(x)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in(1,+\infty).\end{cases}$
Hence we have
$\|\Delta\xi_{k}\|_{p}^{p}$ =$=\int_{\{x\in\mathbb{R}^{n}\colon~{}1<|x|<2\}}|\Delta\eta_{k}|^{p}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big{)}^{p}r^{n-1}\,\mathrm{d}r$
$\quad+\omega_{n-1}\int_{k^{-\frac{1}{n}}}^{1}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}r^{-2}\big{)}^{p}r^{n-1}\,\mathrm{d}r$
$=O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{p}+\frac{\omega_{n-1}}{n-2p}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{p}\bigg{(}1-\frac{1}{k^{\frac{n-2p}{n}}}\bigg{)}$
$\leq O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{p}+\frac{\omega_{n-1}}{n-2p}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{p}$
$=O((\ln k)^{-\frac{2p}{n}})+O((\ln k)^{-\frac{2p}{n}})~{}\frac{1}{k^{\frac{n-2p}{n}}}$
and
$\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}$ =$=\int_{\{x\in\mathbb{R}^{n}\colon~{}1<|x|<2\}}|\Delta\eta_{k}|^{\frac{n}{2}}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big{)}^{\frac{n}{2}}r^{n-1}\,\mathrm{d}r$
$\quad+\omega_{n-1}\int_{k^{-\frac{1}{n}}}^{1}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}r^{-2}\big{)}^{\frac{n}{2}}r^{n-1}\,\mathrm{d}r$
$=O((\ln k)^{-1})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{\frac{n}{2}}+\omega_{n-1}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{\frac{n}{2}}\frac{\ln k}{n}$
$=O((\ln k)^{-1})+\frac{\omega_{n-1}}{n}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}\big{)}^{\frac{n}{2}}=O((\ln k)^{-1})+1.$
This yields at once that
$\|\xi_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}=\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}+\|\Delta\xi_{k}\|_{p}^{\frac{n}{2}}\to 1\quad\text{as}~{}k\to\infty.$
It follows that $\|\xi_{k}\|\to 1$ as $k\to\infty$. Let us assume $\alpha>\beta_{\gamma,n}$, then there exists $\delta>0$ such that $\alpha=(1+\delta)\beta_{\gamma,n}$. Consequently, one has
$\sup_{u\in E,~{}\|u\|\leq 1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$$\geq\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<k^{-\frac{1}{n}}\}}\frac{\Phi_{\alpha,j_{0}}(\xi_{k}/\|\xi_{k}\|)}{|x|^{\gamma}}\,\mathrm{d}x$
$=\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\frac{\Phi_{\alpha,j_{0}}\bigg{(}\frac{1}{\|\xi_{k}\|}\bigg{(}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}r^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\bigg{)}\bigg{)}}{r^{\gamma-n+1}}\,\mathrm{d}r$
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\Phi_{\alpha,j_{0}}\bigg{(}\frac{1}{\|\xi_{k}\|}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{)}$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\operatorname{\exp}\bigg{(}\alpha\bigg{|}\frac{1}{\|\xi_{k}\|}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{|}^{\frac{n}{n-2}}\bigg{)}$
$=\frac{\omega_{n-1}}{(n-\gamma)}~{}\operatorname{\exp}\bigg{[}\bigg{(}1-\frac{\gamma}{n}\bigg{)}\ln k\bigg{(}\frac{1+\delta}{\|\xi_{k}\|^{\frac{n}{n-2}}}-1\bigg{)}\bigg{]}\to+\infty\quad\text{as}\quad k\to\infty,$
which completes the proof of Theorem 1.3. ∎
As a byproduct of the above theorem, we have the following corollary.
Corollary 3.1.
Suppose that $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Then for all $\alpha\geq 0$ and $u\in E$, there holds
$\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\in L^{1}(\mathbb{R}^{n}).$
Proof.
Since $C^{\infty}_{0}(\mathbb{R}^{n})$ in dense in $E$, for a given $\varepsilon>0$, one can assume $\psi\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u-\psi\|<\varepsilon$. Define
$S_{1}:=\{x\in\mathbb{R}^{n}\colon~{}|u(x)-\psi(x)|\leq 1\}\quad\text{and}\quad S_{2}:=\{x\in\mathbb{R}^{n}\colon~{}|u(x)-\psi(x)|>1\}.$
Notice that $|u(x)|\leq 1+\sup_{x\in\mathbb{R}^{n}}|\psi(x)|:=C$ for all $x\in S_{1}$, where $C>0$ is a suitable constant. It follows that $\Phi_{\alpha,j_{0}}(u)\lesssim|u|^{\frac{nj_{0}}{n-2}}$, where $\frac{nj_{0}}{n-2}\geq p^{\ast}$. Hence by using Lemma 2.5, we have
$\int_{S_{1}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\int_{S_{1}}\frac{|u|^{\frac{nj_{0}}{n-2}}}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\int_{\{x\in S_{1}:~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in S_{1}:~{}|x|\geq 1\}}|u|^{\frac{nj_{0}}{n-2}}\,\mathrm{d}x<+\infty.$ (3.5)
Recall that for a given $\eta>0$ and $\wp>1$, there holds
$(a+b)^{\wp}\leq(1+\eta)a^{\wp}+C_{\eta}b^{\wp},\quad\forall~{}a,b>0,$ (3.6)
where $C_{\eta}=(1-(1+\eta)^{-\frac{1}{\wp-1}})^{1-\wp}$. Moreover, by using (3.6) and Young’s inequality, we have
$\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\int_{S_{2}}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}}+C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\textit{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x.$
Let $\eta,\varepsilon>0$ be sufficiently small enough such that $2\alpha(1+\eta)\varepsilon^{\frac{n}{n-2}}<\beta_{\gamma,n}$ holds. Hence we obtain by using the definition of $S_{2}$ and Theorem 1.3 that
$\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\alpha(1+\eta),j_{0}}(u-\psi)}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u-\psi/\|u-\psi\|)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
In addition, since $\psi$ has compact support, we have
$\int_{\textit{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
Thus it follows that
$\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$ (3.7)
In view of (3.5) and (3.7), we conclude that the statement of the corollary is well established. ∎
The proof of Theorem 1.4 is based on the following crucial lemma. More precisely, we have
Lemma 3.2.
Let the hypotheses of Theorem 1.4 be satisfied. Then there exists a sequence $\{\overset{\sim}{u}_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|\overset{\sim}{u}_{k}\|=1$ and $\overset{\sim}{u}_{k}\rightharpoonup u$ in $E$ as $k\to\infty$. In addition, for some suitable conatant $C$ independent of $k$, there holds
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,\quad\forall~{}\ell^{\prime}\in(\ell,L_{n}(u)).$
Proof.
By the standard density argument, we can assume a sequence $\{\psi_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u_{k}-\psi_{k}\|<\frac{1}{k},~{}\forall~{}k\in\mathbb{N}$. Suppose that $\overset{\sim}{u}_{k}:=\frac{\psi_{k}}{\|\psi_{k}\|},~{}\forall~{}k\in\mathbb{N}$. It follows that the sequence $\{\overset{\sim}{u}_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ and satisfying
$\|\overset{\sim}{u}_{k}\|=1,\quad\|u_{k}-\overset{\sim}{u}_{k}\|\leq\frac{2}{k},~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad\overset{\sim}{u}_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty.$ (3.8)
Divide the whole space $\mathbb{R}^{n}$ as follows
$S_{1}:=\{x\in\mathbb{R}^{n}\colon~{}|u_{k}(x)|\leq 2\}\quad\text{and}\quad S_{2}:=\{x\in\mathbb{R}^{n}\colon~{}|u_{k}(x)|>2\}.$
Notice that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\int_{S_{1}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{S_{2}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x:=J_{1}+J_{2}.$
Due to Lemma 2.5, we obtain for all $k\in\mathbb{N}$ that
$J_{1}$ $=\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big{|}\frac{u_{k}}{2}\Big{|}^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big{|}\frac{u_{k}}{2}\Big{|}^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|<1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|\geq 1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|\geq 1\}}|u_{k}|^{p^{\ast}}\,\mathrm{d}x\leq C(\gamma,\ell,n),$ (3.9)
where $C(\gamma,\ell,n)>0$ is a constant depending only on $\gamma,\ell$ and $n$ but independent on $k$. Further, we can also write
$J_{2}=\int_{\Gamma_{1}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\Gamma_{2}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x:=J_{2,\Gamma_{1}}+J_{2,\Gamma_{2}},$
where
$\Gamma_{1}:=\{x\in S_{2}\colon~{}|u_{k}(x)-\overset{\sim}{u}_{k}(x)|\leq 1\}\quad\text{and}\quad\Gamma_{2}:=\{x\in S_{2}\colon~{}|u_{k}(x)-\overset{\sim}{u}_{k}(x)|>1\}.$
By direct calculation, one has $|\overset{\sim}{u}_{k}(x)|>1$ for all $x\in\Gamma_{1}$ and $k\in\mathbb{N}$. In virtue of (3.6) with $\eta=\frac{\ell^{\prime}-\ell}{\ell}$, we infer that
$J_{2,\Gamma_{1}}$ $\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell^{\prime}\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.10)
We again split the integral $J_{2,\Gamma_{2}}$ as follows
$J_{2,\Gamma_{2}}=\int_{\Gamma_{21}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\Gamma_{22}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x:=J_{2,\Gamma_{21}}+J_{2,\Gamma_{22}},$
where
$\Gamma_{21}:=\Gamma_{2}\cap\{x\in\mathbb{R}^{n}\colon~{}|\overset{\sim}{u}_{k}(x)|<1\}\quad\text{and}\quad\Gamma_{22}:=\Gamma_{2}\cap\{x\in\mathbb{R}^{n}\colon~{}|\overset{\sim}{u}_{k}(x)|\geq 1\}.$
Utilizing the inequality (3.6) with $\eta=1$, we obtain from the definition of $\Gamma_{2}$ that
$J_{2,\Gamma_{21}}$ $\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(2\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{1}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.11)
Set $w_{k}:=\frac{u_{k}-\overset{\sim}{u}_{k}}{\|u_{k}-\overset{\sim}{u}_{k}\|}$ and take $k_{0}\in\mathbb{N}$ sufficiently large enough such that $2\ell\beta_{\gamma,n}\big{(}\frac{2}{k}\big{)}^{\frac{n}{n-2}}<\beta_{\gamma,n}$ for all $k\geq k_{0}$. Then by using (3.8), Corollary 2.6 and Theorem 1.3, we obtain for all $k\geq k_{0}$ that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(w_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq\overset{\sim}{C},$
for some suitable constant $\overset{\sim}{C}>0$, which is independent on the choice of $k$. In addition, by using Corollary 3.1, one has
$C=\max\bigg{\{}\overset{\sim}{C},\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{1}-\overset{\sim}{u}_{1})}{|x|^{\gamma}}\,\mathrm{d}x,\dots,\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k_{0}}-\overset{\sim}{u}_{k_{0}})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{\}}<+\infty.$
In conclusion, we deduce that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~{}\forall~{}k\in\mathbb{N}.$ (3.12)
From (3) and (3.12), one can notice that
$J_{2,\Gamma_{21}}=\int_{\Gamma_{21}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~{}\forall~{}k\in\mathbb{N},$ (3.13)
for some suitable constant $C>0$ independent on the choice of $k$. Let $t,t^{\prime}>1$ be such $\frac{1}{t}+\frac{1}{t^{\prime}}=1$, where $t=\frac{2\ell^{\prime}}{\ell+\ell^{\prime}}$. Choose $\eta=\frac{\ell^{\prime}-\ell}{2\ell}$, then by using (3.6) and Young’s inequality, we have
$J_{2,\Gamma_{22}}$ $\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}((1+\eta)\ell\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\frac{1}{t}\int_{\Gamma_{22}}\frac{\operatorname{\exp}(\ell^{\prime}\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x+\frac{1}{t^{\prime}}\int_{\Gamma_{22}}\frac{\operatorname{\exp}(t^{\prime}C_{\eta}\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq C_{1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{2}\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,$ (3.14)
where $C_{1},~{}C_{2}>0$ are some suitable constants independent on the choice of $k$. Let $k_{0}\in\mathbb{N}$ be sufficiently large enough such that $t^{\prime}C_{\eta}\ell\beta_{\gamma,n}\big{(}\frac{2}{k}\big{)}^{\frac{n}{n-2}}<\beta_{\gamma,n}$ for all $k\geq k_{0}$. Then arguing as before, it is not difficult to prove that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C_{3},~{}\forall~{}k\in\mathbb{N},$ (3.15)
for some suitable constant $C_{3}>0$ independent of $k$. Thus by using (3) and (3.15), we infer that
$\sup_{k\in\mathbb{N}}J_{2,\Gamma_{22}}\leq C_{1}\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{4},$ (3.16)
where $C_{4}>0$ is some suitable constant independent on the choice of $k$. Hence by using (3), (3), (3.13) and (3.16), we conclude the proof. ∎
Proof of Theorem 1.4..
The proof of the first part of Theorem 1.4 is based on method of contradiction. Due to Lemma 3.2, it suffices to prove Theorem 1.4 for compactly supported smooth functions. Indeed, let $\{u_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ be such that for any $\ell\in(0,L_{n}(u))$, we have
$\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\neq 0\quad\text{in}\quad E\quad\text{as}\quad k\to\infty\quad\text{and}\quad\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x=+\infty.$ (3.17)
Let $u_{k}^{\ast}$ be the Schwarz symmetrization of $u_{k}$. Notice that $(1/|x|^{\gamma})^{\ast}=1/|x|^{\gamma}$ and therefore by applying Hardy-Littlewood inequality, for instance, see Lemma 2.1 and (2.2), we have
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x\leq\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.18)
From (3.17) and (3.18), one has
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow\infty\quad\text{as}\quad k\to\infty.$ (3.19)
Let $R>0$ be fixed and join us in writing
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\int_{B_{R}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,\quad\forall~{}k\in\mathbb{N}.$
It follows directly that
$\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq\frac{1}{R^{\gamma}}\int_{B_{R}^{c}}\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})\,\ \mathrm{d}x.$
Due to the monotone convergence theorem together with Lemma 2.5 and (3.17) yields
$\int_{B_{R}^{c}}\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})\,\mathrm{d}x$ $=\int_{B_{R}^{c}}\sum_{j=j_{0}}^{\infty}\frac{(\ell\beta_{\gamma,n})^{j}}{j!}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x=\sum_{j=j_{0}}^{\infty}\int_{B_{R}^{c}}\frac{(\ell\beta_{\gamma,n})^{j}}{j!}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x$
$\leq\frac{(\ell\beta_{\gamma,n})^{j_{0}}}{j_{0}!}\|u_{k}\|^{\frac{nj_{0}}{n-2}}_{\frac{nj_{0}}{n-2}}+\sum_{j=j_{0}+1}^{\infty}\frac{(\ell\beta_{\gamma,n})^{j}}{j!}\int_{B_{R}^{c}}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x$
$\leq\frac{(\ell\beta_{\gamma,n})^{j_{0}}}{j_{0}!}c_{1}+\sum_{j=j_{0}+1}^{\infty}\frac{(\ell\beta_{\gamma,n})^{j}}{j!}\int_{B_{R}^{c}}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x,$
where $c_{1}>0$ is a suitable constant independent on the choice of $k$. Now by using Lemma 2.4, Lemma 2.5 and (3.17), one sees that
$|u_{k}^{\ast}(x)|\leq\frac{1}{|x|^{\frac{n}{p^{\ast}}}}\Big{(}\frac{n}{\omega_{n-1}}\Big{)}^{\frac{1}{p^{\ast}}}\|u_{k}\|_{p^{\ast}}\leq\Big{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\Big{)}^{\frac{1}{p^{\ast}}}\frac{1}{|x|^{\frac{n}{p^{\ast}}}},~{}\forall~{}~{}x\neq 0,$
where $c_{2}>0$ is again a suitable constant independent on the choice of $k$. The above inequality with the fact that $\frac{n^{2}j}{(n-2)p^{\ast}}>n$ for all $j\geq j_{0}+1$ implies that
$\int_{B_{R}^{c}}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x$ $\leq\omega_{n-1}\Big{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\Big{)}^{\frac{nj}{(n-2)p^{\ast}}}\int_{R}^{\infty}r^{n-\frac{n}{p^{\ast}}\frac{nj}{n-2}-1}\,\mathrm{d}r$
$\leq\frac{\omega_{n-1}(n-2)p^{\ast}R^{n}}{n\big{(}n(j_{0}+1)-(n-2)p^{\ast}\big{)}}\bigg{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}R^{n}}\bigg{)}^{\frac{nj}{(n-2)p^{\ast}}},~{}\forall~{}j\geq j_{0}+1.$
On combining all the above information, one has that
$\int_{B_{R}^{c}}\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})\,\mathrm{d}x$ $\leq\frac{(\ell\beta_{\gamma,n})^{j_{0}}}{j_{0}!}c_{1}+\frac{\omega_{n-1}(n-2)p^{\ast}R^{n}}{n\big{(}n(j_{0}+1)-(n-2)p^{\ast}\big{)}}\operatorname{\exp}\Bigg{[}\ell\beta_{\gamma,n}\bigg{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}R^{n}}\bigg{)}^{\frac{n}{(n-2)p^{\ast}}}\Bigg{]}$
$:=C(\ell,\gamma,R,n)<+\infty,$
where $C(\ell,\gamma,R,n)>0$ is a constant depending only on $\ell,\gamma,R$ and $n$. In conclusion, we infer that
$\sup_{k\in\mathbb{N}}\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C(\ell,\gamma,R,n)<+\infty.$ (3.20)
In addition, a direct consequence of (3.19) and (3.20) implies that
$\lim_{k\to\infty}\int_{B_{R}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}^{\ast}(x)|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x=\infty.$ (3.21)
Further, assume that $g_{k},g\in C^{\infty}_{0}(\mathbb{R}^{n})$ for all $k\in\mathbb{N}$ and consider the following problems
$-\Delta u_{k}=g_{k}\quad\text{in}\quad\mathbb{R}^{n}\quad\text{and}\quad-\Delta u=g\quad\text{in}\quad\mathbb{R}^{n}.$
It is obvious that $g_{k},g\in L^{t}(\mathbb{R}^{n})$ for all $k\in\mathbb{N}$ and for any $t\in\{p,\frac{n}{2}\}$. Define $g_{k}^{\ast}(x)=g_{k}^{\sharp}(\sigma_{n}|x|^{n})$ for all $x\in\mathbb{R}^{n}$ and $k\in\mathbb{N}$. Hence we obtain from (3.17) that $\|g_{k}\|_{t}\leq 1$ for all $k\in\mathbb{N}$ and $g_{k}\rightharpoonup g$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$. Now by using Lemma 2.2, there exists a subsequence of $\{g_{k}\}_{k}$ still not relabelled and $h\in L^{t}([0,+\infty))$ such that
$g_{k}^{\sharp}\to h~{}\text{a.e. in}~{}[0,+\infty)~{}~{}\text{as}~{}~{}k\to\infty~{}~{}\text{and}~{}~{}\|h\|_{L^{t}([0,+\infty))}\geq\|g^{\sharp}\|_{L^{t}([0,+\infty))}=\|g\|_{t}~{}~{}~{}\text{for any }~{}t\in\{p,\frac{n}{2}\}.$ (3.22)
It follows from Lemma 2.3, (2.1) and integration by parts formula that, for all $0<s_{1}<s_{2}<+\infty$, one has
$\quad\quad u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})$ $\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+D(n)\|g_{k}\|_{\frac{n}{2}}\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\bigg{(}\int_{0}^{t}g_{k}^{\sharp}(\eta)\,\ \mathrm{d}\eta\bigg{)}t^{\frac{2}{n}-2}\,\mathrm{d}t+D(n)$
$=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\Bigg{[}\frac{1}{s_{1}^{1-\frac{2}{n}}}\int_{0}^{s_{1}}g_{k}^{\sharp}(t)\,\mathrm{d}t-\frac{1}{s_{2}^{1-\frac{2}{n}}}\int_{0}^{s_{2}}g_{k}^{\sharp}(t)\,\mathrm{d}t+\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t\Bigg{]}+D(n),~{}\forall~{}k\in\mathbb{N}.$
Due to Hölder’s inequality and (3.22), we obtain for all $s\in(0,+\infty)$ that
$\frac{1}{s^{1-\frac{2}{n}}}\int_{0}^{s}g_{k}^{\sharp}(t)\,\mathrm{d}t\leq\|g_{k}^{\sharp}\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}=\|g_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\leq 1,~{}\forall~{}k\in\mathbb{N}.$
Moreover, we deduce from the above inequalities that there exists a suitable constant $\overset{\sim}{D}(n)>0$ which is depending on $n$ but independent on the choice of $k$ and there holds
$u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\ \mathrm{d}t+\overset{\sim}{D}(n),~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad 0<s_{1}<s_{2}<+\infty.$ (3.23)
Define the function $w_{k}$ on $(0,|B_{R}|]$ by
$w_{k}(s):=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s}^{|B_{R}|}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\ \mathrm{d}t,~{}\forall~{}s\in(0,|B_{R}|].$
Notice that the map $\cdot\mapsto w_{k}(\cdot)$ is decreasing on $(0,|B_{R}|]$ and $w_{k}(|B_{R}|)=0$. It follows from (3.17), (3.23), Lemma 2.4 and Lemma 2.5 that
$u_{k}^{\sharp}(s)$ $\leq w_{k}(s)+u_{k}^{\sharp}(|B_{R}|)+\overset{\sim}{D}(n)\leq w_{k}(s)+\overset{\sim}{D}(n,R),~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|),$
[…]
$|u_{k}^{\sharp}(s)|^{\frac{n}{n-2}}\leq(1+\eta)|w_{k}(s)|^{\frac{n}{n-2}}+C_{\eta}\overset{\sim}{D}(n,R)^{\frac{n}{n-2}},~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|).$
This together with the properties of the Schwarz symmetrization ensures that
$\int_{B_{R}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}^{\ast}(x)|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}^{\sharp}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s$
$\leq\operatorname{exp}(\ell\beta_{\gamma,n}C_{\eta}\overset{\sim}{D}(n,R)^{\frac{n}{n-2}})\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s.$
It follows from the above inequality and (3.21) that
$\lim_{k\to\infty}\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s=\infty.$ (3.24)
In virtue of Hölder’s inequality and (3.22), for all $s\in(0,|B_{R}|)$, we have
$w_{k}(s)$ $\leq\bigg{(}\int_{0}^{|B_{R}|}|g_{k}^{\sharp}(t)|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\int_{s}^{|B_{R}|}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg{)}^{1-\frac{2}{n}}$
$\leq\|g^{\ast}_{k}\|_{\frac{n}{2}}\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}}\leq\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}},~{}\forall~{}k\in\mathbb{N}.$ (3.25)
Next, we claim that for any (¯$\ell^{\prime}\in(\bar{\ell},L_{n}(u),~{}k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$, there exist $k\in\mathbb{N}$ and $s\in(0,s_{0})$ satisfying
$w_{k}(s)\geq\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}},~{}\forall~{}k>k_{0}.$
Indeed, if not, let there exists $k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$ be such that for some (¯$\ell^{\prime}\in(\bar{\ell},L_{n}(u)$, we have
$w_{k}(s)<\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}},~{}\forall~{}k\geq k_{0}\quad\text{and}\quad s\in(0,s_{0}).$ (3.26)
By using (3) and (3.26), for sufficiently large $k$, we have
$\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})$ $\,\mathrm{d}s=\bigg{(}\int_{0}^{s_{0}}+\int_{s_{0}}^{|B_{R}|}\bigg{)}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s$
$\leq\int_{0}^{s_{0}}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\bigg{(}\frac{|B_{R}|}{s}\bigg{)}^{(1-\frac{\gamma}{n})\frac{\bar{\ell}}{\ell^{\prime}}}\,\mathrm{d}s+\int_{s_{0}}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\bigg{(}\frac{|B_{R}|}{s}\bigg{)}^{(1-\frac{\gamma}{n})\bar{\ell}}\,\mathrm{d}s$
<+<$\infty$.< + ∞ .
In view of the above inequality and (3.24), we arrive at a contradiction. This completes the proof of our claim. Thus there exists a subsequence $\{s_{k}\}_{k}\subset(0,s_{0})$ with $s_{k}\leq\frac{1}{k}$ for all $k\in\mathbb{N}$ such that we can assume
$w_{k}(s_{k})\geq\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}.$ (3.27)
For $L>0$ and $v\geq 0$, we define
$H^{L}(v):=\min\{v,L\}~{}\quad\text{and}\quad H_{L}(v):=v-H^{L}(v).$
In view of (3.27), one can easily see that $w_{k}(s_{k})\nearrow\infty$ as $k\to\infty$ and also we have $w_{k}(|B_{R}|)=0$. This shows that we can find a subsequence of $\{s_{k}\}_{k}$ still not relabelled and the existence of $\xi_{k}\in(s_{k},|B_{R}|)$ satisfying $w_{k}(\xi_{k})=L$ for each $k\in\mathbb{N}$. In addition, since the map $\cdot\mapsto g_{k}^{\sharp}(\cdot)$ is always nonnegative and nonincreasing, therefore by using Hölder’s inequality, one has
$\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\xi_{k}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t\leq\frac{ng_{k}^{\sharp}(s_{k})}{2\beta(n,2)^{1-\frac{2}{n}}}\Big{(}|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big{)}.$
The above inequality together with (3.27) implies that
$\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}-L\leq w_{k}(s_{k})-w_{k}(\xi_{k})\leq\frac{ng_{k}^{\sharp}(s_{k})}{2\beta(n,2)^{1-\frac{2}{n}}}\Big{(}|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big{)},~{}\forall~{}k\in\mathbb{N}.$ (3.28)
Now from (3.28), we infer that $g_{k}^{\sharp}(s_{k})\nearrow\infty$ as $k\to\infty$ and thus there exists $k_{0}\in\mathbb{N}$ such that $g_{k}^{\sharp}(s_{k})>L$ for all $k\geq k_{0}$. Consequently, up to a subsequence of $\{s_{k}\}_{k}$ still not relabelled, we can find $\eta_{k}\in(s_{k},|B_{R}|)$ satisfying $g_{k}^{\sharp}(\eta_{k})=L$ for each $k\in\mathbb{N}$ and $g_{k}^{\sharp}(\eta)<L$ for all $\eta>\eta_{k}$. For all $k\geq k_{0}$, we set $\theta_{k}:=\min\{\eta_{k},\xi_{k}\}$, then one sees from (3.27) that
$\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}-L$ $\leq w_{k}(s_{k})-w_{k}(\xi_{k})=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\xi_{k}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$
$=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\Bigg{[}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+\int_{\theta_{k}}^{\xi_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+L\int_{s_{k}}^{\xi_{k}}\frac{1}{t^{1-\frac{2}{n}}}\,\mathrm{d}t\Bigg{]}$
$\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\Bigg{[}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+\frac{nL}{2}\Big{(}|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big{)}\Bigg{]}.$ (3.29)
On the other hand, the application of Hölder’s inequality yields
$\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$ $\leq\bigg{(}\int_{s_{k}}^{\theta_{k}}|g_{k}^{\sharp}(t)-L|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\int_{s_{k}}^{\theta_{k}}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg{)}^{1-\frac{2}{n}}$
$\leq\bigg{(}\int_{0}^{\eta_{k}}|g_{k}^{\sharp}(t)-L|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\int_{s_{k}}^{|B_{R}|}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg{)}^{1-\frac{2}{n}}$
$\leq\bigg{(}\int_{0}^{|B_{R}|}|H_{L}(g_{k}^{\sharp}(t))|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}$
$\leq\|H_{L}(g^{\sharp}_{k})\|_{L^{\frac{n}{2}}([0,+\infty))}\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}.$ (3.30)
If possible, let us choose $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$. It is obvious that $\ln\big{(}\frac{|B_{R}|}{s_{k}}\big{)}\to\infty$ as $k\to\infty$. Hence we obtain from (3) and (3) as $k\to\infty$ that
$\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}$ $\leq\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+o_{k}(1)$
$\leq\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$ (3.31)
It may be deduced from direct calculations that
11 11 $=\|u_{k}\|^{\frac{n}{2}}=\|g_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g_{k}\|^{\frac{n}{2}}_{p}$
$=\|g^{\ast}_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g^{\ast}_{k}\|^{\frac{n}{2}}_{p}=\|g^{\sharp}_{k}\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}$
$\geq\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}$
$\quad+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$ (3.32)
In view of (3) and (3), we get as $k\to\infty$ that
1($1-\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}\geq\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$ (3.33)
By the definition of $H_{L}$ and (3.22), for a.a. $s\in[0,+\infty)$, one has
$h(s)-H_{L}(h(s))\to h(s)\quad\text{as}\quad L\to\infty\quad\text{and}\quad g^{\sharp}_{k}(s)-H_{L}(g^{\sharp}_{k}(s))\to h(s)-H_{L}(h(s))\quad\text{as}\quad k\to\infty.$
Letting $k\to\infty$ in (3.33) and applying Fatou’s lemma, we have
1($1-\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}\geq\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$
Despite this, again by Fatou’s lemma, we obtain from the above inequality as $L\to\infty$ that
$1-\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}$ $\geq\|h\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|h\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}\geq\|g^{\sharp}\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}$
$=\|g^{\ast}\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g^{\ast}\|^{\frac{n}{2}}_{p}=\|g\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g\|^{\frac{n}{2}}_{p}=\|u\|^{\frac{n}{2}}=1-\bigg{(}\frac{1}{L_{n}(u)}\bigg{)}^{\frac{n-2}{2}},$
which contradicts $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$ and thus we complete the proof of the first part of Theorem 1.4.
To complete the proof, we only have to prove the sharpness of $L_{n}(u)$. For this, we shall have to contruct a sequence $\{u_{k}\}_{k}\subset E$ and a function $u\in E$ satisfying
$\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty,\quad\|u\|=\delta<1$ (3.34)
but
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow\infty\quad\text{as}\quad k\to\infty,\quad\forall~{}\ell>L_{n}(u).$
Let $\{\xi_{k}\}_{k}\subset E$ be a sequence of radial functions, as defined in (3.4). It follows from the standard arguments that
$\xi_{k}\rightharpoonup 0\quad\text{in}\quad E,\quad\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\to 1\quad\text{and}\quad\|\Delta\xi_{k}\|^{p}_{p}\to 0\quad\text{as}\quad k\to\infty.$
Take a radial function $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u\|=\delta<1$ and $\text{supp}(u)\subset\{x\in\mathbb{R}^{n}:~{}2<|x|<3\}$. Define $v_{k}:=u+(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\xi_{k}$ for all $k\in\mathbb{N}$. Notice that $\Delta u$ and $\Delta\xi_{k}$ have disjoint supports. It follows at once that
$\|\Delta v_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}=\|\Delta u\|_{\frac{n}{2}}^{\frac{n}{2}}+(1-\delta^{\frac{n}{2}})+\zeta_{k}\quad\text{and}\quad\|\Delta v_{k}\|^{p}_{p}=\|\Delta u\|^{p}_{p}+\zeta_{k},$
where $\zeta_{k}\to 0$ as $k\to\infty$. Consequently, one sees that $\|v_{k}\|=1+\zeta_{k}$. Denote $u_{k}:=\frac{v_{k}}{1+\zeta_{k}}$. This implies the validity of (3.34). Moreover, for any $\ell>L_{n}(u)$, there exists $\varepsilon>0$ such that we can assume $\ell=(1+\varepsilon)L_{n}(u)$. In conclusion, we have the following esimates
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}$ $\,\mathrm{d}x\geq\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<k^{-\frac{1}{n}}\}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}\big{(}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\xi_{k}\big{)}}{|x|^{\gamma}}\,\mathrm{d}x$
$=\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}\bigg{(}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\bigg{(}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}r^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\bigg{)}\bigg{)}}{r^{\gamma-n+1}}\,\mathrm{d}r$
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\Phi_{\ell\beta_{\gamma,n},j_{0}}\bigg{(}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{)}$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\operatorname{\exp}\bigg{(}\ell\beta_{\gamma,n}\bigg{|}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{|}^{\frac{n}{n-2}}\bigg{)}$
$=\frac{\omega_{n-1}}{(n-\gamma)}~{}\operatorname{\exp}\bigg{[}\bigg{(}1-\frac{\gamma}{n}\bigg{)}\ln k\bigg{(}\frac{1+\varepsilon}{(1+\zeta_{k})^{\frac{n}{n-2}}}-1\bigg{)}\bigg{]}\to+\infty\quad\text{as}\quad k\to\infty,$
from which the proof of Theorem 1.4 is well established. ∎
4. Proof of Theorem 1.5 and Theorem 1.6
Basically, the purpose of this section is to prove Theorem 1.5 and Theorem 1.6. More precisely, we shall demonstrate the asymptotic behavior for the supremum $\text{ATSC}(\ell,\gamma)$ for the subcritical sharp singular Adams’ type inequality. In fact, we also mention here that the proof of Theorem 1.5 doesn’t require the validity of $\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma)<+\infty$. Additionally, we also provide the relationship between the suprema for the critical and subcritical sharp singular Adams’ type inequalities. Now we establish the following crucial lemmas, which are useful in the proof of Theorem 1.5 and Theorem 1.6.
Lemma 4.1.
There holds
$\text{ATSC}(\ell,\gamma)=\sup_{u\in E,~{}\|\Delta u\|_{\frac{n}{2}}\leq 1,~{}\|\Delta u\|_{p}=1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x.$
Proof.
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$. Define
$w(x)=u(\lambda x)\quad\text{with}\quad\lambda=\|\Delta u\|_{p}^{\frac{p}{n-2p}}>0.$
By direct calculations, one has
$\Delta w(x)=\lambda^{2}\Delta u(\lambda x),\quad\|\Delta w\|_{\frac{n}{2}}=\|\Delta u\|_{\frac{n}{2}}\quad\text{and}\quad\|\Delta w\|_{p}=\lambda^{-\frac{n-2p}{p}}\|\Delta u\|_{p}=1.$
It follows that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(w)}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u(\lambda x))}{|x|^{\gamma}}\,\mathrm{d}x=\frac{1}{\lambda^{n-\gamma}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$
$=\cfrac{1}{\|\Delta u\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x,$
from which we conclude the proof. ∎
Lemma 4.2.
The subcritical sharp singular Adams’ type inequality in $E$ is a consequence of the critical sharp singular Adams’ type inequality in $E$. More specifically, if $\text{ATC}_{a,b}(\gamma)<+\infty$, then $\text{ATSC}(\ell,\gamma)<+\infty$. Furthermore, there holds
$\text{ATSC}(\ell,\gamma)\leq\Bigg{(}\cfrac{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATC}_{a,b}(\gamma).$
In particular, if $\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma)<+\infty$, then
$\text{ATSC}(\ell,\gamma)\leq\Bigg{(}\cfrac{\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}}\Bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma).$
Proof.
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Define
$w(x)=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{n}}u(\lambda x)\quad\text{with}\quad\lambda=\Bigg{(}\cfrac{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}\Bigg{)}^{\frac{p}{b(n-2p)}}.$
By direct calculations, we have
$\|\Delta w\|^{a}_{\frac{n}{2}}=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\big{(}\frac{n-2}{n}\big{)}a}\|\Delta u\|^{a}_{\frac{n}{2}}\leq\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\big{(}\frac{n-2}{n}\big{)}a}$
and
$\|\Delta w\|^{b}_{p}$ $=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{(\frac{n-2}{n})b}\lambda^{-b\big{(}\frac{n-2p}{p}\big{)}}\|\Delta u\|^{b}_{p}=1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{(\frac{n-2}{n})a}.$
It follows that $\|\Delta w\|^{a}_{\frac{n}{2}}+\|\Delta w\|^{b}_{p}\leq 1$. Consequently, we deduce from the definition of $\text{ATC}_{a,b}(\gamma)$ that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u(\lambda x))}{|\lambda x|^{\gamma}}\,\mathrm{d}(\lambda x)=\lambda^{n-\gamma}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(w)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\Bigg{(}\cfrac{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATC}_{a,b}(\gamma),$
from which the proof of Lemma 4.2 is well established. ∎
Note that Lemma 4.1 allows us to assume $\|\Delta u\|_{p}=1$ in the subcritical sharp singular Adams’ type inequality. Now we can prove Theorem 1.5.
Proof of Theorem 1.5..
On exploiting the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $E$ and Lemma 4.1, without loss of generality, one can assume $u\in C_{0}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$, $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Further, since $|u|=\sqrt{u^{2}}$, we can easily deduce that $|\Delta|u||=|\Delta u|$. Let us define
$S:=\Bigg{\{}x\in\mathbb{R}^{n}\colon~{}|u(x)|>\bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n}}\bigg{\}}.$
Now using again the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $D^{2,p}(\mathbb{R}^{n})$ and the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$, one has that $S$ is a bounded domain. Indeed, we have
$|S|=\int_{S}1\,\mathrm{d}x$ $\leq\cfrac{1}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}}}~{}\|u\|^{p^{\ast}}_{p^{\ast}}\leq\cfrac{C(n,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}}},$
where $C(n,p)>0$ is a constant depending only on $n$ and $p$. Now join us in writing
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x=\int_{S}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\mathbb{R}^{n}\setminus S}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x=:I_{1}+I_{2}.$
To evaluate $I_{2}$, we first notice that $\mathbb{R}^{n}\setminus S\subset\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}$. By using the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ and the fact that $\|\Delta u\|_{p}=1$, we have
$I_{2}$$\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big{(}\ell(1-\frac{\gamma}{n})\big{)}^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big{(}\ell(1-\frac{\gamma}{n})\big{)}^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1,~{}|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1,~{}|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\leq C(n,\gamma,p),$ (4.1)
where $C(n,\gamma,p)>0$ is a constant depending only on $n$, $\gamma$ and $p$. Finally, to evaluate $I_{1}$, we define
$w(x):=|u(x)|-\bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2}{n}}\quad\text{in}\quad S.$
It follows that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and $\|\Delta w\|_{\frac{n}{2}}\leq 1$. By employing (3.6) with $\eta=\frac{\beta(n,2)}{\ell}-1$, we have
$|u|^{\frac{n}{n-2}}$ $\leq\Bigg{(}|w|+\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n}}\Bigg{)}^{\frac{n}{n-2}}\leq(1+\eta)|w|^{\frac{n}{n-2}}+C_{\eta}\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{-\frac{2}{n-2}}\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+1.$
Gathering all the above information, we obtain from the singular Adams’ inequality [32, Theorem 1.2] with homogeneous Navier boundary condition that
$I_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}(\ell(1-\frac{\gamma}{n})|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{S}\frac{\operatorname{\exp}(\beta_{\gamma,n}|w|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$ (4.2)
where $C(n,\gamma,p)>0$ is a constant depending only on $n$, $\gamma$ and $p$. Hence we obtain from (4) and (4) that
$\text{ATSC}(\ell,\gamma)\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}.$
To prove the sharpness of $\beta(n,2)$, that is, $\text{ATSC}(\beta(n,2),\gamma)=+\infty$, inspired by Adams [2], we consider $f\in C^{\infty}([0,1])$ satisfying
$f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0;\quad f(1)=f^{\prime}(1)=1\quad\text{and}\quad f^{\prime\prime}(1)=0.$
Fix $0<\varepsilon<\frac{1}{2}$ and define the function
$\Psi(s)=\begin{cases}\varepsilon f(\frac{s}{\varepsilon})\quad\quad\quad\quad\quad\quad\text{if}\quad 0<s\leq\varepsilon,\\ s\quad\quad\quad\quad\quad\quad\quad\quad\text{if}\quad\varepsilon<s\leq 1-\varepsilon,\\ 1-\varepsilon f(\frac{1-s}{\varepsilon})\quad\quad\quad\text{if}\quad 1-\varepsilon<s\leq 1,\\ 1\quad\quad\quad\quad\quad\quad\quad\quad\text{if}\quad s>1,\end{cases}$
with $\text{supp}(\Psi)\subset(0,+\infty)$. Let $v\colon~{}(0,+\infty)\to\mathbb{R}$ be such that
$v(s)=\Psi\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)},~{}\forall~{}s>0\quad\text{and}\quad r\in(0,1).$ (4.3)
It follows that $v(|x|)\in E$ and $v(|x|)=1$ for all $x\in B_{r}$. Moreover, one can easily obtain
$\Delta v(|x|)=(-1)\frac{\alpha(n,2)}{\omega_{n-1}}\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}\Big{(}\log\frac{1}{r}\Big{)}^{-1}|x|^{-2}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-2}\Big{)}|x|^{-2},$
where $\alpha(n,2)=\omega_{n-1}(n-2)$ and the big $O$ term involves terms that are constant mutiple of $\Psi^{\prime\prime}\Big{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\Big{)}\big{(}\log\frac{1}{r}\big{)}^{-2}$. In addition, by direct computations, we also have
$\int_{\mathbb{R}^{n}}|\Delta v(|x|)|^{\frac{n}{2}}\,\mathrm{d}x=\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Big{(}\log\frac{1}{r}\Big{)}^{1-\frac{n}{2}}\Theta_{\frac{n}{2}},$
where
$\Theta_{\frac{n}{2}}$ $=\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{0}^{\infty}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}$
$=\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Bigg{(}\int_{0}^{r}+\int_{r}^{r^{1-\varepsilon}}+\int_{r^{1-\varepsilon}}^{r^{\varepsilon}}+\int_{r^{\varepsilon}}^{1}+\int_{1}^{\infty}\Bigg{)}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}$
$\leq 1+2\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}},$
where we used the following estimates
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{r^{1-\varepsilon}}^{r^{\varepsilon}}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq 1;$
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{a}^{b}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}=0,~{}\forall~{}(a,b)\in\{(0,r),(1,+\infty)\}$
and
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{a}^{b}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq$ $~{}\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}},$
$~{}\forall~{}(a,b)\in\{(r,r^{1-\varepsilon}),(r^{\varepsilon},1)\}.$
Similarly, we can prove that
$\int_{\mathbb{R}^{n}}|\Delta v(|x|)|^{p}\,\mathrm{d}x=\omega_{n-1}^{1-p}\alpha(n,2)^{p}\Big{(}\log\frac{1}{r}\Big{)}^{1-p}\Theta_{p},$
where
$\Theta_{p}\leq 1+2\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{p}.$
Define $\bar{v}\colon(0,+\infty)\to\mathbb{R}$ by
$\bar{v}(s)=\Big{(}\log\frac{1}{r}\Big{)}^{1-\frac{2}{n}}v(s),~{}\forall~{}s>0\quad\text{and}\quad r\in(0,1).$
It follows immediately that
$\bar{v}(|x|)=\Big{(}\log\frac{1}{r}\Big{)}^{1-\frac{2}{n}},~{}\forall~{}x\in B_{r},\quad\|\Delta\bar{v}\|^{\frac{n}{2}}_{\frac{n}{2}}=\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}$
and
$\|\Delta\bar{v}\|^{p}_{p}=\Big{(}\log\frac{1}{r}\Big{)}^{1-\frac{2p}{n}}\omega_{n-1}^{1-p}\alpha(n,2)^{p}\Theta_{p}<\Big{(}\frac{1}{r}\Big{)}^{1-\frac{2p}{n}}\omega_{n-1}^{1-p}\alpha(n,2)^{p}\Theta_{p},$
where the last inequality is obtained by using the elementary inequality $\frac{x-1}{x}\leq\log x\leq x-1,~{}\forall~{}x>0$. In virtue of the above information, one has
$\text{ATSC}((\beta(n,2),\gamma)$ $\geq\cfrac{1}{\Big{\|}\Delta\Big{(}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big{)}\Big{\|}_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<r\}}\cfrac{\Phi_{\beta_{\gamma,n},j_{0}}\Big{(}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big{)}}{|x|^{\gamma}}\,\mathrm{d}x$
$>\frac{\omega_{n-1}}{n-\gamma}~{}C(n,\varepsilon,r,\|f^{\prime}\|_{\infty})~{}r^{(1+n)(1-\frac{\gamma}{n})}~{}\Phi_{\beta_{\gamma,n},j_{0}}\Bigg{(}\cfrac{\big{(}\log\frac{1}{r}\big{)}^{1-\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}\Bigg{)}$
$\gtrsim\frac{\omega_{n-1}}{n-\gamma}~{}C(n,\varepsilon,r,\|f^{\prime}\|_{\infty})~{}r^{1-\frac{\gamma}{n}}~{}\Phi_{\beta_{\gamma,n},j_{0}}\Bigg{(}\cfrac{\big{(}\log\frac{1}{r}\big{)}^{1-\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}\Bigg{)}$
$\sim\frac{\omega_{n-1}}{n-\gamma}~{}C(n,\varepsilon,r,\|f^{\prime}\|_{\infty})~{}r^{1-\frac{\gamma}{n}}~{}\operatorname{\exp}\bigg{(}\beta_{\gamma,n}\bigg{|}\cfrac{\big{(}\log\frac{1}{r}\big{)}^{1-\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}\bigg{|}^{\frac{n}{n-2}}\bigg{)}$
$=\frac{\omega_{n-1}}{n-\gamma}~{}C(n,\varepsilon,r,\|f^{\prime}\|_{\infty})\operatorname{\exp}\Bigg{[}\bigg{(}1-\frac{\gamma}{n}\bigg{)}\ln\frac{1}{r}\Bigg{(}\frac{n}{\Big{(}1+2\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}}\Big{)}^{\frac{2}{n-2}}}-1\Bigg{)}\Bigg{]}$
$\to+\infty\quad\text{as}\quad r\to 0^{+},~{}\text{if we take }~{}\varepsilon>0~{}\text{sufficiently small enough},$
where
$C(n,\varepsilon,r,\|f^{\prime}\|_{\infty})=\Bigg{(}\cfrac{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-p}\alpha(n,2)^{p}\Theta_{p}\big{)}^{\frac{1}{p}}}\Bigg{)}^{p^{\ast}(1-\frac{\gamma}{n})}.$
This shows the validity of the sharpness of $\beta(n,2)$ and hence the claim is well established. Next, we consider the same sequence $\{\xi_{k}\}_{k}\subset E$ that is used in the proof of Theorem 1.3. Observe that the following estimates hold
$\|\Delta\xi_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}=1+O((\ln k)^{-1})\quad\text{and}\quad\|\Delta\xi_{k}\|^{p}_{p}\leq C_{1}(\ln k)^{-\frac{2p}{n}}+C_{2}(\ln k)^{-\frac{2p}{n}}\frac{1}{k^{\frac{n-2p}{n}}},$
where $C_{1}$ and $C_{2}$ are two positive constants.
Define
$w_{k}=\frac{\xi_{k}}{\|\Delta\xi_{k}\|_{\frac{n}{2}}}.$
It follows at once that
$\|\Delta w_{k}\|_{\frac{n}{2}}=1\quad\text{and}\quad\|\Delta w_{k}\|^{p}_{p}\leq\|\Delta\xi_{k}\|^{p}_{p}\leq C_{1}(\ln k)^{-\frac{2p}{n}}+C_{2}(\ln k)^{-\frac{2p}{n}}\frac{1}{k^{\frac{n-2p}{n}}}.$
Moreover, we have
$\text{ATSC}(\ell,\gamma)$ $\geq\cfrac{1}{\|\Delta w_{k}\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<k^{-\frac{1}{n}}\}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(w_{k})}{|x|^{\gamma}}\,\mathrm{d}x$
$\geq\frac{\omega_{n-1}(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}{\big{(}C_{1}+C_{2}k^{-(\frac{n-2p}{n})}\big{)}^{\frac{p^{\ast}}{p}(1-\frac{\gamma}{n})}}\int_{0}^{k^{-\frac{1}{n}}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}\bigg{(}\frac{1}{\|\Delta\xi_{k}\|_{\frac{n}{2}}}\bigg{(}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}r^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\bigg{)}\bigg{)}}{r^{\gamma-n+1}}\,\mathrm{d}r$
$\geq\frac{(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}{\big{(}C_{1}+C_{2}k^{-(\frac{n-2p}{n})}\big{)}^{\frac{p^{\ast}}{p}(1-\frac{\gamma}{n})}}~{}\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}\bigg{(}\frac{1}{\|\Delta\xi_{k}\|_{\frac{n}{2}}}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{)}$
$\sim\frac{(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}{\big{(}C_{1}+C_{2}k^{-(\frac{n-2p}{n})}\big{)}^{\frac{p^{\ast}}{p}(1-\frac{\gamma}{n})}}~{}\frac{\omega_{n-1}}{(n-\gamma)}~{}\operatorname{\exp}\Bigg{[}\Big{(}1-\frac{\gamma}{n}\Big{)}\frac{\ell}{\beta(n,2)}\ln k\Bigg{(}\frac{1}{\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{n-2}}}-\frac{\beta(n,2)}{\ell}\Bigg{)}\Bigg{]}.$
For sufficiently large $k$, one has
$\frac{1}{\big{(}C_{1}+C_{2}k^{-(\frac{n-2p}{n})}\big{)}^{\frac{p^{\ast}}{p}(1-\frac{\gamma}{n})}}\thickapprox\frac{1}{C_{1}^{\frac{p^{\ast}}{p}(1-\frac{\gamma}{n})}}\quad\text{and}\quad\frac{1}{\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{n-2}}}-\frac{\beta(n,2)}{\ell}\thickapprox 1-\frac{\beta(n,2)}{\ell}.$
In conclusion, we have
$\text{ATSC}(\ell,\gamma)\gtrsim(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}~{}\operatorname{\exp}\Big{[}\Big{(}1-\frac{\gamma}{n}\Big{)}\ln k\Big{(}\frac{\ell}{\beta(n,2)}-1\Big{)}\Big{]}.$
Notice that when $\ell\underset{\approx}{<}\beta(n,2)$, we can able to choose $k$ sufficiently large enough such that
$\ln k\thickapprox\frac{p}{1-\frac{\ell}{\beta(n,2)}}\quad\text{and }\quad\Big{(}1-\frac{\gamma}{n}\Big{)}\ln k\Big{(}\frac{\ell}{\beta(n,2)}-1\Big{)}\thickapprox-\Big{(}1-\frac{\gamma}{n}\Big{)}p.$
Therefore, whenever $\ell\underset{\approx}{<}\beta(n,2)$, we have
$\text{ATSC}(\ell,\gamma)\gtrsim\cfrac{\operatorname{\exp}\Big{(}-\Big{(}1-\frac{\gamma}{n}\Big{)}p\Big{)}}{\Big{(}1-\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}\thickapprox\cfrac{\frac{n-2}{2}~{}\operatorname{\exp}\Big{(}-\Big{(}1-\frac{\gamma}{n}\Big{)}p\Big{)}}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$
where we used
$\frac{1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}}{1-\frac{\ell}{\beta(n,2)}}\thickapprox\frac{n-2}{2}\quad\text{for}\quad\ell\underset{\approx}{<}\beta(n,2).$
This yields that
$\text{ATSC}(\ell,\gamma)\geq\cfrac{c(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$
where $c(n,\gamma,p)>0$ is a constant depending only on $n$, $\gamma$ and $p$. This finishes the proof. ∎
Proof of Theorem 1.6..
Let $0<b\leq\frac{n}{2}$ and $u\in E\setminus\{0\}$ be such that $\|\Delta u\|^{a}_{\frac{n}{2}}+\|\Delta u\|^{b}_{p}\leq 1$. Define
$\|\Delta u\|_{\frac{n}{2}}=\theta\in(0,1)\quad\text{and}\quad\|\Delta u\|^{b}_{p}\leq 1-\theta^{a}.$
If $\frac{1}{2^{\frac{n-2p}{p}}}<\theta<1$, then we define
$w(x)=\frac{u(\lambda x)}{\theta}\quad\text{and}\quad\lambda=\bigg{(}\frac{(1-\theta^{a})^{\frac{1}{b}}}{\theta}\bigg{)}^{\frac{p}{n-2p}}.$
This yields that
$\|\Delta w\|_{\frac{n}{2}}=1\quad\text{and}\quad\|\Delta w\|^{p}_{p}=\frac{\|\Delta u\|^{p}_{p}}{\lambda^{n-2p}\theta^{p}}\leq\frac{(1-\theta^{a})^{\frac{p}{b}}}{\lambda^{n-2p}\theta^{p}}=1.$
Now by employing Theorem 1.5, one has
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u(\lambda x))}{|\lambda x|^{\gamma}}\,\mathrm{d}(\lambda x)\leq\lambda^{n-\gamma}\int_{\mathbb{R}^{n}}\frac{\Phi_{\theta^{\frac{n}{n-2}}\beta_{\gamma,n},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\lambda^{n-\gamma}\text{ATSC}(\theta^{\frac{n}{n-2}}\beta(n,2),\gamma)\leq\bigg{(}\frac{(1-\theta^{a})^{\frac{1}{b}}}{\theta}\bigg{)}^{p^{\ast}(1-\frac{\gamma}{n})}\frac{C(n,\gamma,p)}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}$
$\leq\frac{(1-\theta^{a})^{\frac{p^{\ast}}{b}(1-\frac{\gamma}{n})}}{(1-\theta^{\frac{n}{2}})^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}C(n,\gamma,p)\leq C(n,\gamma,p,a,b),$
thanks to the fact that $b\leq\frac{n}{2}$, where $C(n,\gamma,p,a,b)>0$ is a constant depending only on $n,\gamma,p,a$ and $b$. If $0<\theta\leq\frac{1}{2^{\frac{n-2p}{p}}}$, then we consider
$w(x)=2^{\frac{n-2p}{p}}u(2x).$
It follows directly that
$\|\Delta w\|_{\frac{n}{2}}=2^{\frac{n-2p}{p}}\|\Delta u\|_{\frac{n}{2}}\leq 1\quad\text{and}\quad\|\Delta w\|_{p}=\|\Delta u\|_{p}\leq 1.$
Due to Theorem 1.5, we have
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u(2x))}{|2x|^{\gamma}}\,\mathrm{d}(2x)\leq 2^{n-\gamma}\int_{\mathbb{R}^{n}}\frac{\Phi_{2^{-\big{(}\frac{n^{2}}{(n-2)p^{\ast}}\big{)}}\beta_{\gamma,n},j_{0}}(w)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq 2^{n-\gamma}\text{ATSC}\Big{(}2^{-\big{(}\frac{n^{2}}{(n-2)p^{\ast}}\big{)}}\beta(n,2),\gamma\Big{)}$
$\leq 2^{n-\gamma}\frac{C(n,\gamma,p)}{\Big{(}1-2^{-\frac{n^{2}}{2p^{\ast}}}\Big{)}^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}}\leq C(n,\gamma,p),$
where $C(n,\gamma,p)>0$ is a constant depending only on $n,~{}\gamma$ and $p$. In conclusion, we deduce that $\text{ATC}_{a,b}(\gamma)<+\infty$ for $b\leq\frac{n}{2}$. Next, to prove the sharpness of $\beta(n,2)$, inspired by Theorem 1.5, we assume that
$\bar{v}(s)=\Big{(}\log\frac{1}{r}\Big{)}^{1-\frac{2}{n}}v(s),~{}\forall~{}s>0\quad\text{and}\quad r\in(0,1),$
where $v$ is defined as in (4.3). Let us choose
$\bar{w}(s)=\delta_{r}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}},~{}\forall~{}s>0\quad\text{and}\quad\delta_{r}\in(0,1),$
with $\delta_{r}\to 1$ as $r\to 0^{+}$ and there holds
$\delta^{a}_{r}+\delta^{b}_{r}\frac{\|\Delta\bar{v}\|^{b}_{p}}{\|\Delta\bar{v}\|^{b}_{\frac{n}{2}}}=1.$
Notice that $\|\Delta\bar{w}\|^{a}_{\frac{n}{2}}+\|\Delta\bar{w}\|^{b}_{p}=1$. Hence, by choosing $\ell>\beta(n,2)$, we obtain
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(\bar{w})}{|x|^{\gamma}}\,\mathrm{d}x$$\geq\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<r\}}\cfrac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}\Big{(}\delta_{r}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big{)}}{|x|^{\gamma}}\,\mathrm{d}x$
$=\frac{\omega_{n-1}}{n-\gamma}~{}r^{n-\gamma}~{}\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}\bigg{(}\frac{\delta_{r}(\log\frac{1}{r})^{1-\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}\bigg{)}$
$\sim\frac{\omega_{n-1}}{n-\gamma}~{}r^{n-\gamma}~{}\operatorname{\exp}\Bigg{[}\ell\Big{(}1-\frac{\gamma}{n}\Big{)}\Bigg{|}\frac{\delta_{r}(\log\frac{1}{r})^{1-\frac{2}{n}}}{\big{(}\omega_{n-1}^{1-\frac{n}{2}}\alpha(n,2)^{\frac{n}{2}}\Theta_{\frac{n}{2}}\big{)}^{\frac{2}{n}}}\Bigg{|}^{\frac{n}{n-2}}\Bigg{]}$
$=\frac{\omega_{n-1}}{n-\gamma}~{}\operatorname{\exp}\Bigg{[}n(1-\frac{\gamma}{n}\Big{)}\log\frac{1}{r}\Bigg{(}\frac{\ell}{\beta(n,2)}\frac{\delta_{r}^{\frac{n}{n-2}}}{\Big{(}1+2\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}}\Big{)}^{\frac{2}{n-2}}}-1\Bigg{)}\Bigg{]}$
$\to+\infty\quad\text{as}\quad r\to 0^{+},~{}\text{if we take }~{}\varepsilon>0~{}\text{sufficiently small enough}.$
This finishes the proof of the sharpness of $\beta(n,2)$. Finally, we deduce from Lemma 4.2 that
$\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)\leq\text{ATC}_{a,b}(\gamma).$ (4.4)
Let $\{v_{k}\}_{k}\subset E\setminus\{0\}$ be a maximizing sequence of $\text{ATC}_{a,b}(\gamma)$. It follows that $\|\Delta v_{k}\|^{a}_{\frac{n}{2}}+\|\Delta v_{k}\|^{b}_{p}\leq 1$ and there holds
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(v_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\text{ATC}_{a,b}(\gamma).$
Define
$w_{k}(x)=\frac{v_{k}(\lambda_{k}x)}{\|\Delta v_{k}\|_{\frac{n}{2}}},\quad\text{where}\quad\lambda_{k}=\bigg{(}\frac{1-\|\Delta v_{k}\|^{a}_{\frac{n}{2}}}{\|\Delta v_{k}\|^{b}_{\frac{n}{2}}}\bigg{)}^{\frac{p^{\ast}}{nb}}.$
Thus, we have
$\|\Delta w_{k}\|_{\frac{n}{2}}=1\quad\text{and}\quad\|\Delta w_{k}\|_{p}=\frac{\|\Delta v_{k}\|_{p}}{\lambda_{k}^{\frac{n-2p}{p}}\|\Delta v_{k}\|_{\frac{n}{2}}}\leq\frac{(1-\|\Delta v_{k}\|^{a}_{\frac{n}{2}})^{\frac{1}{b}}}{\lambda_{k}^{\frac{n-2p}{p}}\|\Delta v_{k}\|_{\frac{n}{2}}}=1.$
Consequently, we obtain from the definition of $\text{ATSC}(\ell,\gamma)$ that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(v_{k})}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(v_{k}(\lambda_{k}x))}{|\lambda_{k}x|^{\gamma}}\,\mathrm{d}(\lambda_{k}x)\leq\lambda_{k}^{n-\gamma}\int_{\mathbb{R}^{n}}\frac{\Phi_{\|\Delta v_{k}\|_{\frac{n}{2}}^{\frac{n}{n-2}}\beta_{\gamma,n},j_{0}}(w_{k})}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\lambda_{k}^{n-\gamma}\text{ATSC}\Big{(}\|\Delta v_{k}\|_{\frac{n}{2}}^{\frac{n}{n-2}}\beta(n,2),\gamma\Big{)}$
$\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$
From the above inequality, we have
$\text{ATC}_{a,b}(\gamma)\leq\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$ (4.5)
In view of (4.4) and (4.5), we deduce that (1.15) holds.
To complete the proof, we claim that $\text{ATC}_{a,b}(\gamma)<+\infty$ implies $b\leq\frac{n}{2}$. Indeed, if not, let there exists some $b>\frac{n}{2}$ such that $\text{ATC}_{a,b}(\gamma)<+\infty$. Thus, one can obtain from (1.15) that
$\limsup_{\ell\to\beta(n,2)^{-}}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)<+\infty.$ (4.6)
On the other hand, by Theorem 1.5, we have
$\limsup_{\ell\to\beta(n,2)^{-}}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma)>0.$ (4.7)
In virtue of (4.6) and (4.7), one has
[…]
which is a contradiction, thanks to $b>\frac{n}{2}$ and hence the claim. This finishes the proof. ∎
5. Proof of Theorem 1.7
This section concerns about the attainability of the best constant in the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. For this purpose, we first need to prove Theorem 1.7.
Proof of Theorem 1.7..
First, we claim that the continuous embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ holds for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$. Indeed, let $v\in E$ and $\eta,~{}\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$ satisfying $\gamma\eta^{\prime}<n$. By using Hölder’s inequality and Lemma 2.5, one has
$\int_{\mathbb{R}^{n}}\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{B_{1}}\frac{|v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B^{c}_{1}}\frac{|v|^{\varrho}}{|x|^{\gamma}}~{}\mathrm{d}x$
$\leq\bigg{(}\int_{B_{1}}|v|^{\varrho\eta}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}\bigg{(}\int_{B_{1}}\frac{1}{|x|^{\gamma\eta^{\prime}}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta^{\prime}}}+\int_{\mathbb{R}^{n}}|v|^{\varrho}\,\mathrm{d}x\lesssim\|v\|^{\varrho},$
which conclude the claim. It remains to check the compactness of the above embedding. To prove this, take a bounded sequence $\{v_{k}\}_{k}\subset E$, then one can find a subsequence $\{v_{k_{j}}\}_{j}$ and $v$ in $E$ satisfying
$v_{k_{j}}\rightharpoonup v~{}\text{in}~{}E,~{}v_{k_{j}}\to v~{}\text{in}~{}L^{\theta}(B_{R})\quad\text{for any}~{}R>0~{}\text{and}~{}\theta\geq 1,~{}v_{k_{j}}\to v~{}\text{a.e. in}~{}\mathbb{R}^{n}~{}\text{as}~{}j\to\infty,$
we thank to Lemma 2.5. Due to the Egoroff’s theorem, we can notice that for any open ball $B_{R}$ and $\delta>0$, there exists $A_{\delta}\subset B_{R}$ with $|A_{\delta}|<\delta$ satisfying $v_{k_{j}}\to v$ uniformly in $B_{R}\setminus A_{\delta}$ as $j\to\infty$. Further, let us write
$\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$
$=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x+\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{B_{R}\setminus A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$
$\quad+\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x:=J_{1}+J_{2}+J_{3}.$ (5.1)
Suppose that $1<t<\frac{n}{\gamma}$ and $1<r<\frac{n}{t\gamma}$ such that $\frac{1}{\tau}+\frac{1}{\tau^{\prime}}=1$ for any $\tau\in\{t,r\}$. Now by using Hölder’s inequality and Lemma 2.5, we have
$\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq|A_{\delta}|^{\frac{1}{t^{\prime}}}\bigg{(}\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho t}}{|x|^{t\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{t}}\leq|A_{\delta}|^{\frac{1}{t^{\prime}}}\bigg{(}\int_{A_{\delta}}\frac{1}{|x|^{rt\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{rt}}\bigg{(}\int_{A_{\delta}}|v_{k_{j}}-v|^{\varrho r^{\prime}t}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}t}}$
$\leq|A_{\delta}|^{\frac{1}{t^{\prime}}}\bigg{(}\int_{B_{R}}\frac{1}{|x|^{rt\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{rt}}\bigg{(}\int_{\mathbb{R}^{n}}|v_{k_{j}}-v|^{\varrho r^{\prime}t}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}t}}$
$\lesssim|A_{\delta}|^{\frac{1}{t^{\prime}}}\bigg{(}\int_{B_{R}}\frac{1}{|x|^{rt\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{rt}}\sup_{j\in\mathbb{N}}\|v_{k_{j}}-v\|^{\varrho}\lesssim|A_{\delta}|^{\frac{1}{t^{\prime}}},$
which directly implies that
$J_{1}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.2)
Further, we also have
$\int_{B_{R}\setminus A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x\leq\sup_{B_{R}\setminus A_{\delta}}|v_{k_{j}}-v|^{\varrho}\int_{B_{R}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\sup_{B_{R}\setminus A_{\delta}}|v_{k_{j}}-v|^{\varrho},$
which together with the fact that $v_{k_{j}}\to v$ uniformly in $B_{R}\setminus A_{\delta}$ as $j\to\infty$ ensures that
$J_{2}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{B_{R}\setminus A_{\delta}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.3)
Finally, one can observe that
$\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x\leq\frac{1}{R^{\gamma}}\int_{\mathbb{R}^{n}}|v_{k_{j}}-v|^{\varrho}\,\mathrm{d}x\lesssim\frac{1}{R^{\gamma}}\sup_{j\in\mathbb{N}}\|v_{k_{j}}-v\|^{\varrho}\lesssim\frac{1}{R^{\gamma}},$
which implies at once that
$J_{3}=\lim_{R\to\infty}\lim_{\delta\to 0}\lim_{j\to\infty}\int_{\mathbb{R}^{n}\setminus B_{R}}\frac{|v_{k_{j}}-v|^{\varrho}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (5.4)
In virtue of (5.2), (5.3) and (5.4), we obtain from (5) that $v_{k_{j}}\to v$ in $L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ as $j\to\infty$. This completes the proof. ∎
The following corollary is an immediate consequence of the above theorem.
Corollary 5.1.
The best constant $S_{\varrho}$ in the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$ is achieved, where $S_{\varrho}$ is given by
$S_{\varrho}:=\inf_{u\in E\setminus\{0\}}\frac{\|u\|}{~{}\|u\|_{\varrho,\gamma}}.$
Proof.
Indeed, let $\{v_{k}\}_{k}\subset E$ be a sequence such that $\|v_{k}\|_{\varrho,\gamma}=1~{}\text{and}~{}\|v_{k}\|\to S_{\varrho}~{}\text{ as}~{}k\to\infty$. It follows that $\{v_{k}\}_{k}\subset E$ is bounded. Thus, up to a subsequence, still denoted by itself and $v\in E$ such that
$v_{k}\rightharpoonup v~{}\text{in}~{}E,~{}v_{k}\to v~{}\text{in}~{}L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x),~{}\forall~{}\varrho\geq p^{\ast}~{}\text{and}~{}v_{k}\to v~{}\text{a.e. in}~{}\mathbb{R}^{n}~{}\text{as}~{}k\to\infty,$
thanks to Theorem 1.7. This infers that $\|v\|_{\varrho,\gamma}=\lim_{k\to\infty}\|v_{k}\|_{\varrho,\gamma}=1$. Moreover, due to the weakly lower semicontinuity of the norm, we have $\|v\|\leq\liminf_{k\to\infty}\|v_{k}\|=S_{\varrho}\leq\|v\|$. It follows that the best constant $S_{\varrho}$ is achieved. This finishes the proof. ∎
6. Proof of Theorem 1.8
In this section, we shall use the mountain pass theorem, standard topological tools, singular Adams’ inequality, a new concentration-compactness principle due to Lions (see, for example, Theorem 1.3 and Theorem 1.4 ) and Theorem 1.7, respectively, to prove Theorem 1.8. Evidently, the problem ($\boldsymbol{S}$) has a variational structure. Indeed, since we are interested in studying ($\boldsymbol{S}$), we define the Euler–Lagrange variational energy functional $J\colon E\to\mathbb{R}$ by
$J(u):=\frac{1}{p}\|\Delta u\|^{p}_{p}+\frac{2}{n}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x,~{}\forall~{}u\in E.$ (6.1)
It is easy to see that $J$ is well-defined, of class $C^{1}(E,\mathbb{R})$ and its Gâteaux derivative is given by
$\langle{J^{\prime}(u),v}\rangle:=\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta v\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta v\,\mathrm{d}x-\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}v\,\mathrm{d}x,~{}\forall~{}u,v\in E.$
Moreover, the critical points of $J$ are exactly weak solutions of ($\boldsymbol{S}$). To prove our existence results, we need a version of the mountain pass theorem of Rabinowitz [42], as stated below.
Definition 6.1.
(Palais-Smale compactness condition) Let $X$ be a Banach space and $\mathcal{I}:X\to\mathbb{R}$ be a functional of class $C^{1}(X,\mathbb{R})$. We say that $\mathcal{I}$ satisfies the Palais-Smale compactness condition at a suitable level $c\in\mathbb{R}$, if for any sequence $\{u_{k}\}_{k}\subset X$ such that
$\mathcal{I}(u_{k})\to c\quad\text{and}\quad\sup_{\|\varphi\|_{X}=1}|\langle{\mathcal{I}^{\prime}(u_{k}),\varphi}\rangle|\to 0\quad\text{as}\quad k\to\infty$ (6.2)
has a strongly convergent subsequence in $X$. Note that the sequence $\{u_{k}\}_{k}\subset X$ satisfying (6.2) is known as a Palais-Smale sequence at level $c\in\mathbb{R}$, which is denoted by the symbol $(PS)_{c}$.
Theorem 6.2.
(The mountain pass theorem ) Let $X$ be a real Banach space and $I\in C^{1}(X,\mathbb{R})$. Suppose $I(0)=0$ and there hold
  • (a)
    there exist two constants $\alpha,\rho>0$ such that $I(u)\geq\alpha$ for all $u\in X$ with $\|u\|=\rho$;
  • (b)
    there exists $e\in X$ satisfying $\|e\|>\rho$ such that $I(e)<0$.
Define
$\Gamma:=\{\gamma\in C^{1}([0,1],X)\colon\gamma(0)=0,~{}\gamma(1)=e\}.$
Then
$c:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}I(\gamma(t))\geq\alpha$
and there exists a $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ for $I$ in $X$. Consequently, if $I$ satisfies the Palais-Smale condition, then $I$ possesses a critical value $c\geq\alpha$.
To use Theorem 6.2, we first prove the mountain pass geometry of the energy functional $J$ as follows.
Lemma 6.3.
Let the hypotheses $(g_{1})-(g_{4})$ be satisfied. Then the following results are true:
  • (i)
    there exist $\delta,\rho>0$ such that $J(u)\geq\delta$ for all $u\in E$ and $\|u\|=\rho$;
  • (ii)
    there exists $e\in E$ with $\|e\|>\rho$ such that $J(e)<0$.
Proof.
Let $t,t^{\prime}>1$ be such that $\frac{1}{t}+\frac{1}{t^{\prime}}=1$. Further, we assume $\kappa>0$ such that $\alpha t^{\prime}\kappa^{\frac{n}{n-2}}\leq\beta_{\gamma,n}$ whenever $\|u\|\leq\kappa$. From (1.2), we obtain by using Hölder’s inequality, Theorem 1.7 and Corollary 2.6 that
$\int_{\mathbb{R}^{n}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\zeta\|u\|^{\tau}_{\tau,\gamma}+D_{\zeta}\|u\|^{q}_{qt,\gamma}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha t^{\prime},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{t^{\prime}}}$
$\leq\zeta S^{-\tau}_{\tau}\|u\|^{\tau}+D_{\zeta}S^{-q}_{qt}\|u\|^{q}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u/\|u\|)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{t^{\prime}}}$
$\leq C(\zeta\|u\|^{\tau}+\|u\|^{q}),\quad\forall\quad\zeta>0,$
where we used Theorem 1.3 to obtain the constant $C>0$, which is given by
$C:=\max\Bigg{\{}S^{-\tau}_{\tau},D_{\zeta}S^{-q}_{qt}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u/\|u\|)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{t^{\prime}}}\Bigg{\}}<+\infty.$
Define $\rho=\min\{1,\kappa\}$. Then for all $u\in E$ with $\|u\|=\rho$, we have
$J(u)\geq\|u\|^{\frac{n}{2}}\Big{[}\frac{2}{n}-C(\zeta\|u\|^{\tau-\frac{n}{2}}+\|u\|^{q-\frac{n}{2}})\Big{]}.$
Let us assume $0<\zeta<\frac{2}{nC}$. Notice that $\tau,q>\frac{n}{2}$, therefore we may choose $\rho>0$ small enough such that
$\frac{2}{n}-C(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}})>0.$
It follows that
$J(u)\geq\rho^{\frac{n}{2}}\Big{[}\frac{2}{n}-C(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}})\Big{]}:=\delta>0\quad\text{for}\quad\|u\|=\rho.$
This finishes the proof of the first part of Lemma 6.3.
Next, for the proof of the second part, we assume that $u\in E\setminus\{0\}$, $u\geq 0$ with $\|u\|=1$. Define the map $\Psi\colon[1,+\infty)\to\mathbb{R}$ by $\Psi(s):=s^{-\mu}G(x,su)-G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}:=(0,+\infty)$. By using $(g_{4})$, one can easily check that the map $\cdot\mapsto\Psi(\cdot)$ is an increasing function on $[1,+\infty)$ and thus we have $G(x,su)\geq s^{\mu}G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}$. Now choosing $s\geq 1$ sufficiently large enough, one has
$J(su)\leq\frac{2}{p}s^{\frac{n}{2}}-s^{\mu}\int_{\mathbb{R}^{n}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x\to-\infty\quad\text{as}\quad s\to\infty,$
we thank to $\mu>\frac{n}{2}$. The proof is now completed by choosing $e=su$ with $s$ sufficiently large. Hence the lemma is well established. ∎
Due to Lemma 6.3 and Theorem 6.2, we can define the mountain pass minimax level as follows
$c:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}J(\gamma(t))\geq\rho>0,$
where
$\Gamma:=\{\gamma\in C^{1}([0,1],E)\colon\gamma(0)=0,~{}J(\gamma(1))<0\}.$
More precisely, we have the following crucial lemma.
Lemma 6.4.
There exists $\lambda_{0}>0$ sufficiently large enough such that if the hypothesis $(g_{6})$ holds for all $\lambda\geq\lambda_{0}$, then the mountain pass minimax level $c$ satisfies
$0<c<c_{0}:=\min\bigg{\{}\frac{1}{2^{\frac{n}{2}}}\bigg{(}\frac{2\mu-n}{n\mu}\bigg{)}^{\frac{n}{2p}}\bigg{(}\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg{)}^{\frac{n-2}{2}},\frac{1}{2^{p}}\bigg{(}\frac{2\mu-n}{n\mu}\bigg{)}\bigg{(}\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg{)}^{\frac{(n-2)p}{n}}\bigg{\}}.$ (6.3)
Proof.
Let $\psi\in C^{\infty}_{0}(\mathbb{R}^{n},[0,1])$ be a cut-off function such that $\psi\equiv 1$ in $B_{\frac{1}{2}}$, $\psi\equiv 0$ in $B^{c}_{1}$, $|\nabla\psi(x)|\leq 2$ and $|\Delta\psi(x)|\leq 4$ for all $x\in\mathbb{R}^{n}$. By using $(g_{6})$, for any $s\in[0,1]$, we have
$J(s\psi)$ $\leq\frac{s^{p}}{p}\int_{B_{1}}|\Delta\psi|^{p}\,\mathrm{d}x+\frac{2s^{\frac{n}{2}}}{n}\int_{B_{1}}|\Delta\psi|^{\frac{n}{2}}\,\mathrm{d}x-\lambda s^{\vartheta}\int_{B_{\frac{1}{2}}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\bigg{(}\frac{2^{2p}}{p}+\frac{2^{n+1}}{n}\bigg{)}\sigma_{n}s^{p}-\frac{\lambda n\sigma_{n}}{2^{n-\gamma}(n-\gamma)}s^{\vartheta}.$
Define the map $\eta\colon[0,1]\to E$ by $\eta(s)=s\psi$. Further, let us assume $\lambda_{1}>0$ be such that for all $\lambda>\lambda_{1}$, we have
$J(\psi)\leq\bigg{(}\frac{2^{2p}}{p}+\frac{2^{n+1}}{n}\bigg{)}\sigma_{n}-\frac{\lambda_{1}n\sigma_{n}}{2^{n-\gamma}(n-\gamma)}<0.$
It follows that $\eta\in\Gamma$. Moreover, by direct calculations, one has
$c\leq\max_{s\in[0,1]}J(s\psi)$ $\leq\sigma_{n}\max_{s\geq 0}\bigg{\{}\bigg{(}\frac{2^{2p}}{p}+\frac{2^{n+1}}{n}\bigg{)}s^{p}-\frac{\lambda n}{2^{n-\gamma}(n-\gamma)}s^{\vartheta}\bigg{\}}$
$=\sigma_{n}\bigg{(}\frac{\vartheta-p}{\vartheta}\bigg{)}\bigg{(}\frac{2^{2p}}{p}+\frac{2^{n+1}}{n}\bigg{)}^{\frac{\vartheta}{\vartheta-p}}\bigg{(}\frac{2^{n-\gamma}(n-\gamma)p}{\lambda n\vartheta}\bigg{)}^{\frac{p}{\vartheta-p}}\to 0\quad\text{as}\quad\lambda\to\infty.$
This indicates that there exists $\lambda_{0}>\lambda_{1}$ sufficiently large enough such that $c<c_{0}$ for all $\lambda\geq\lambda_{0}$, where $c_{0}$ is defined as in (6.3). This completes the proof. ∎
Lemma 6.5.
Suppose that $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 6.4. Then any $(PS)_{c}$ sequence $\{u_{k}\}_{k}\subset E$ for $J$ is bounded in $E$. Moreover, there holds
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}.$ (6.4)
Proof.
Let $0<c<c_{0}$ and $\{u_{k}\}_{k}\subset E$ be a $(PS)_{c}$ sequence for $J$. It follows that
$J(u_{k})=c+o_{k}(1)\quad\text{and}\quad\bigg{\langle}{J^{\prime}(u_{k}),\frac{u_{k}}{\|u_{k}\|}}\bigg{\rangle}=o_{k}(1)\quad\text{as}\quad k\to\infty.$
By using $(g_{4})$ and the above fact, we have
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|\geq\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\Big{(}\|\Delta u_{k}\|_{p}^{p}+\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\Big{)}\quad\text{as}\quad k\to\infty.$ (6.5)
Indeed, if possible, let $\{u_{k}\}_{k}\subset E$ be unbounded in $E$. Thus we have following situations.
Case-1: Suppose that $\|\Delta u_{k}\|_{p}\to\infty$ and $\|\Delta u_{k}\|_{\frac{n}{2}}\to\infty$ as $k\to\infty$. It follows from $1<p<\frac{n}{2}$ that
$\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\geq\|\Delta u_{k}\|_{\frac{n}{2}}^{p}>1\quad\text{for large}\quad k.$
Thus we obtain from (1.12) and (6.5) that
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|\geq 2^{1-p}\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\|u_{k}\|^{p}\quad\text{as}\quad k\to\infty.$
Dividing $\|u_{k}\|^{p}$ on the both the sides and passing $k\to\infty$, we get $00\geq 2^{1-p}\big{(}\frac{2}{n}-\frac{1}{\mu}\big{)}>$ 0, which is a contradiction.
Case-2: Suppose that $\|\Delta u_{k}\|_{p}\to\infty$ as $k\to\infty$ and $\|\Delta u_{k}\|_{\frac{n}{2}}$ is bounded. It follows from (1.12) and (6.5) that
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|_{E}\geq 2^{1-p}\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\|\Delta u_{k}\|_{p}^{p}\quad\text{as}\quad k\to\infty.$
Dividing $\|\Delta u_{k}\|_{p}^{p}$ on the both the sides and passing $k\to\infty$, we get $00\geq\big{(}\frac{2}{n}-\frac{1}{\mu}\big{)}>$ 0, which is again a contradiction.
Case-3: Suppose that $\|\Delta u_{k}\|_{\frac{n}{2}}\to\infty$ as $k\to\infty$ and $\|\Delta u_{k}\|_{p}$ is bounded, then arguing as before in Case-$2$, we get a contradiction.
From the above three cases, we conclude that $\{u_{k}\}_{k}\subset E$ is a bounded sequence in $E$. Moreover, by using (6.5), one has the following estimates
$\limsup_{k\to\infty}\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\leq\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\quad\text{and}\quad\limsup_{k\to\infty}\|\Delta u_{k}\|_{p}^{\frac{n}{2}}\leq\bigg{(}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\bigg{)}^{\frac{n}{2p}}.$ (6.6)
Recalling the inequality $(a+b)^{\wp}\leq a^{\wp}+b^{\wp}$, $\forall~{}a,b\geq 0$ and $\wp\in(0,1]$, we obtain from (6.6) that
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\bigg{)}^{\frac{2}{n-2}}+\bigg{(}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\bigg{)}^{\frac{n}{(n-2)p}}.$
Notice that $12^{\frac{2}{n-2}},\frac{n\mu}{2\mu-n}>$ 1 and $\frac{2}{n-2}<\frac{n}{(n-2)p}$. Thus one can easily deduce from the above inequality that
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}\bigg{)}^{\frac{2}{n-2}}\Big{(}c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\Big{)}.$ (6.7)
Further, if $c\in(0,1]$, then one has $c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\leq 2c^{\frac{2}{n-2}}$. Therefore, by using (6.7), we infer that
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2^{\frac{n}{2}}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}c\bigg{)}^{\frac{2}{n-2}}.$ (6.8)
On the other hand, if $c\geq 1$, then we have $c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\leq 2c^{\frac{n}{(n-2)p}}$. Now again by using (6.7), we obtain
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2^{\frac{n}{2}}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}c^{\frac{n}{2p}}\bigg{)}^{\frac{2}{n-2}}.$ (6.9)
In view of (6.3), (6.8) and (6.9), we conclude the proof. ∎
Lemma 6.6.
Let the hypotheses $(g_{1})-(g_{4})$ and $(g_{6})$ be satisfied. Let $\{u_{k}\}_{k}\subset E$ be an arbitrary $(PS)_{c}$ sequence for $J$ with $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 6.4. Then there exists a subsequence of $\{u_{k}\}_{k}$ still not relabelled and $u\in E$ such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. In addition, there holds
$|\Delta u_{k}|^{t-2}\Delta u_{k}\rightharpoonup|\Delta u|^{t-2}\Delta u\quad\text{in}\quad L^{\frac{t}{t-1}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty\quad\text{for}\quad t\in\{p,\frac{n}{2}\}.$
Proof.
Let $0<c<c_{0}$ and $\{u_{k}\}_{k}\subset E$ be a $(PS)_{c}$ sequence for $J$. By using Lemma 6.5, $\{u_{k}\}_{k}$ is a bounded sequence in $E$ and (6.4) holds. Hence we can assume without loss of generality, up to a subsequence, still not relabelled and $u\in E$ such that $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$. Now by using Lemma 2.5, Theorem 1.7 and Lemma 2.7, we have
$\begin{cases}u_{k}\to u\quad\text{in}\quad L^{\theta}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for all}\quad\theta\in[1,+\infty);\\ u_{k}\to u\quad\text{in}\quad L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)\quad\text{for all}\quad\varrho\in[p^{\ast},+\infty);\\ \nabla u_{k}\to\nabla u\quad\text{in}\quad L^{t}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for}\quad t\in\{p,\frac{n}{2}\};\\ u_{k}\to u\quad\text{a.e.~{}~{}in}\quad\mathbb{R}^{n}\quad\text{as}\quad k\to\infty.\end{cases}$ (6.10)
Pick $R>0$ and $\eta\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $0\leq\eta\leq 1$ in $\mathbb{R}^{n}$, $\eta\equiv 1$ in $B_{R}$ and $\eta\equiv 0$ in $B^{c}_{2R}$. Moreover, using the fact that $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$ and $J^{\prime}(u_{k})\to 0$ in $E^{\ast}$ as $k\to\infty$, one has
$\langle{J^{\prime}(u_{k})-J^{\prime}(u),(u_{k}-u)\eta}\rangle=o_{k}(1)\quad\text{as}\quad k\to\infty.$ (6.11)
Due to the convexity, we have
$\textit{P}^{t}_{k}(u_{k},u):=(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)(\Delta u_{k}-\Delta u)\geq 0\quad\text{a.e. in}\quad\mathbb{R}^{n}$
for any $k\in\mathbb{N}$ and $t\in\{p,\frac{n}{2}\}$. In view of the well known Simon’s inequality with $n\geq 4$ (for instance, see to (6.27)) and (6.11), there exists $c_{\frac{n}{2}}>0$ such that
$c^{-1}_{\frac{n}{2}}\int_{B_{R}}|\Delta u_{k}-\Delta u|^{\frac{n}{2}}\,\mathrm{d}x$ $\leq\int_{B_{R}}\textit{P}^{\frac{n}{2}}_{k}(u_{k},u)\,\mathrm{d}x\leq\displaystyle\sum_{t\in\{p,\frac{n}{2}\}}\bigg{(}\int_{B_{R}}\textit{P}^{t}_{k}(u_{k},u)\,\mathrm{d}x\bigg{)}$
$\leq\displaystyle\sum_{t\in\{p,\frac{n}{2}\}}\bigg{(}\int_{\mathbb{R}^{n}}\textit{P}^{t}_{k}(u_{k},u)\eta\,\mathrm{d}x\bigg{)}$
$=o_{k}(1)+\int_{\mathbb{R}^{n}}\frac{(g(x,u_{k})-g(x,u))(u_{k}-u)\eta}{|x|^{\gamma}}\,\mathrm{d}x$
$\quad-\sum_{t\in\{p,\frac{n}{2}\}}\Bigg{(}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\Delta\eta(u_{k}-u)\,\mathrm{d}x\Bigg{)}$
$\quad-2\sum_{t\in\{p,\frac{n}{2}\}}\Bigg{(}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x\Bigg{)}\quad\text{as}\quad k\to\infty.$ (6.12)
By employing Hölder’s inequality and (6.10), we obtain for $t\in\{p,\frac{n}{2}\}$ that
$\bigg{|}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-$ $|\Delta u|^{t-2}\Delta u)\Delta\eta(u_{k}-u)\,\mathrm{d}x\bigg{|}$
$\leq\|\Delta\eta\|_{\infty}\big{(}\|\Delta u_{k}\|^{t-1}_{t}+\|\Delta u\|^{t-1}_{t}\big{)}\bigg{(}\int_{B_{2R}}|u_{k}-u|^{t}\,\mathrm{d}x\bigg{)}^{\frac{1}{t}}$
$\lesssim\bigg{(}\int_{B_{2R}}|u_{k}-u|^{t}\,\mathrm{d}x\bigg{)}^{\frac{1}{t}}\to 0\quad\text{as }k\to\infty.$
This yields for $t\in\{p,\frac{n}{2}\}$ that
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\Delta\eta(u_{k}-u)\,\mathrm{d}x=0.$ (6.13)
Similarly, again by using Hölder’s inequality and (6.10), we deduce for $t\in\{p,\frac{n}{2}\}$ that
$\bigg{|}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-$ $|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x\bigg{|}$
$\leq\|\nabla\eta\|_{\infty}\big{(}\|\Delta u_{k}\|^{t-1}_{t}+\|\Delta u\|^{t-1}_{t}\big{)}\bigg{(}\int_{B_{2R}}|\nabla(u_{k}-u)|^{t}\,\mathrm{d}x\bigg{)}^{\frac{1}{t}}$
$\lesssim\bigg{(}\int_{B_{2R}}|\nabla(u_{k}-u)|^{t}\,\mathrm{d}x\bigg{)}^{\frac{1}{t}}\to 0\quad\text{as }k\to\infty.$
It follows that for $t\in\{p,\frac{n}{2}\}$, we have
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x=0.$ (6.14)
Define
$J_{1}:=\zeta\int_{B_{2R}}\frac{(|u_{k}|^{\tau-1}+|u|^{\tau-1})|u_{k}-u|}{|x|^{\gamma}}\,\mathrm{d}x$
and
$J_{2}:=D_{\zeta}\int_{B_{2R}}\frac{\big{(}|u_{k}|^{q-1}\Phi_{\alpha,j_{0}}(u_{k})+|u|^{q-1}\Phi_{\alpha,j_{0}}(u)\big{)}|u_{k}-u|}{|x|^{\gamma}}\,\mathrm{d}x.$
Let $1<s<\frac{n}{\gamma}$ and $r\geq p^{\ast}$ be such that $\frac{1}{\upsilon}+\frac{1}{\upsilon^{\prime}}=1$ for $\upsilon\in\{s,r\}$. Then by using Hölder’s inequality and (6.10), we get
$J_{1}$ $\leq\zeta\Bigg{[}\bigg{(}\int_{B_{2R}}\frac{|u_{k}|^{(\tau-1)r^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}+\bigg{(}\int_{B_{2R}}\frac{|u|^{(\tau-1)r^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}\Bigg{]}\bigg{(}\int_{B_{2R}}\frac{|u_{k}-u|^{r}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\leq\zeta\Big{(}\|u_{k}\|^{\tau-1}_{L^{(\tau-1)r^{\prime}s^{\prime}}(B_{2R})}+\|u\|^{\tau-1}_{L^{(\tau-1)r^{\prime}s^{\prime}}(B_{2R})}\Big{)}\bigg{(}\int_{B_{2R}}\frac{1}{|x|^{\gamma s}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}s}}\bigg{(}\int_{B_{2R}}\frac{|u_{k}-u|^{r}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\lesssim\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{r}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}\to 0\quad\text{as}\quad k\to\infty.$
Due to (6.4), passing if necessary to a subsequence, still denoted by $\{u_{k}\}_{k}$, we can assume
$\sup_{k\in\mathbb{N}}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}.$
Let $\eta,~{}\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$. From the above inequality, we first fix $m\in(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}})$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$ and $\eta>1$ close to $1$ such that $\alpha\eta m<\beta_{\gamma,n}$. It follows from Hölder’s inequality, Theorem 1.3, Corollary 2.6, Corollary 3.1 and (6.10) that
$J_{2}$ $\leq D_{\zeta}\Bigg{[}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{(q-1)\eta^{\prime}}|u_{k}-u|^{\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha\eta,j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}$
$\quad\quad+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u|^{(q-1)\eta^{\prime}}|u_{k}-u|^{\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha\eta,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}\Bigg{]}$
$\lesssim\Bigg{[}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{q\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{q-1}{q\eta^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u_{k}/\|u_{k}\|)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}$
$\quad\quad+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u|^{q\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{q-1}{q\eta^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha\eta,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\eta}}\Bigg{]}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{q\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{q\eta^{\prime}}}$
$\lesssim\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{q\eta^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{q\eta^{\prime}}}\to 0\quad\text{as}\quad k\to\infty.$
In addition, by using (1.1), one sees that
$\bigg{|}\int_{\mathbb{R}^{n}}\cfrac{(g(x,u_{k})-g(x,u))(u_{k}-u)\eta}{|x|^{\gamma}}\,\mathrm{d}x\bigg{|}$ $\leq\int_{B_{2R}}\cfrac{(|g(x,u_{k})|+|g(x,u)|)|u_{n}-u|}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq J_{1}+J_{2}\to 0\quad\text{as}\quad k\to\infty.$
This shows that
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\cfrac{(g(x,u_{k})-g(x,u))(u_{k}-u)\eta}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.15)
Letting $k\to\infty$ in (6) and using (6.13), (6.14) and (6.15), we deduce that $\Delta u_{k}\to\Delta u$ in $L^{\frac{n}{2}}(B_{R})$ as $k\to\infty$ for all $R>0$. Hence, up to a subsequence still denoted by itself such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. It follows that $|\Delta u_{k}|^{t-2}\Delta u_{k}\to|\Delta u|^{t-2}\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$ for $t\in\{p,\frac{n}{2}\}$. Consequently, by using the fact that $\{|\Delta u_{k}|^{t-2}\Delta u_{k}\}_{k}$ is bounded in $L^{\frac{t}{t-1}}(\mathbb{R}^{n})$, we conclude the proof. ∎
Proposition 6.7.
Let the hypotheses $(g_{1})-(g_{6})$ be satisfied. Then the functional $J$ satisfies $(PS)_{c}$ condition for any $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 6.4.
Proof.
Let $\{u_{k}\}_{k}\subset E$ be a $(PS)_{c}$ sequence for $J$ with $0<c<c_{0}$. It follows that
$\frac{1}{p}\|\Delta u_{k}\|^{p}_{p}+\frac{2}{n}\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\to c\quad\text{as}\quad k\to\infty$ (6.16)
and
$|\langle{J^{\prime}(u_{k}),\psi}\rangle|\leq\zeta_{k}\|\psi\|,~{}\forall~{}\psi\in E,$ (6.17)
where $\zeta_{k}\to 0$ as $k\to\infty$. Pick $\psi=u_{k}$ in (6.17), then one has
$\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x-\|\Delta u_{k}\|^{p}_{p}-\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\leq\zeta_{k}\|\ u_{k}\|.$ (6.18)
By using Lemma 6.5, we infer that $\{u_{k}\}_{k}$ is a bounded sequence in $E$ and (6.4) holds. Therefore, one can deduce from (6.16) and (6.18) that
$\int_{\mathbb{R}^{n}}\frac{|G(x,u_{k})|}{|x|^{\gamma}}\,\mathrm{d}x\leq C\quad\text{and}\quad\int_{\mathbb{R}^{n}}\frac{|g(x,u_{k})u_{k}|}{|x|^{\gamma}}\,\mathrm{d}x\leq C$ (6.19)
for some suitable constant $C>0$. Notice that (6.10) also holds. Due to (1.1), Corollary 3.1, Lemma 2.5 and Hölder’s inequality, one has $\frac{g(x,u)}{|x|^{\gamma}}\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})$. Now by employing a similar strategy developed in de Souza-do Ó [19], we have
$\frac{g(x,u_{k})}{|x|^{\gamma}}\to\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$ (6.20)
Now we claim that
$\frac{G(x,u_{k})}{|x|^{\gamma}}\to\frac{G(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$ (6.21)
To prove this, take $R>0$ and let us write
$\int_{\mathbb{R}^{n}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=\int_{B_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x.$
By using $(g_{4})$ and $(g_{5})$, we can find a constant $K_{0}>0$ such that
$\frac{G(x,u_{k})}{|x|^{\gamma}}\leq K_{0}\frac{|g(x,u_{k})|}{|x|^{\gamma}},~{}\forall~{}x\in\mathbb{R}^{n}.$ (6.22)
Hence we obtain by using (6.10), (6.20), (6.22) and the generalized Lebesgue dominated convergence theorem that
$\lim_{R\to\infty}\lim_{k\to\infty}\int_{B_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.23)
Join us in writing
$\int_{B_{R}^{c}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|>K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x$
$\quad\quad+\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|\leq K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x.$
In view of (6.19) and (6.22), we have the following estimates
$\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|>K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\frac{K_{0}}{K}\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|>K\}}\frac{|g(x,u_{k})u_{k}|}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\frac{CK_{0}}{K}+\int_{B_{R}^{c}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x,$
which together with (1.2), Corollary 3.1, Theorem 1.7 and the Lebesgue dominated convergence theorem yields
$\lim_{K\to\infty}\lim_{R\to\infty}\lim_{k\to\infty}\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|>K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$
Due to $(g_{2})$, we can find a constant $C_{1}>0$ such that $|G(x,s)|\leq C_{1}|s|^{\tau}$ for all $|s|\leq K$. It follows from Lemma 2.5 and Theorem 1.7 that
$\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|\leq K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\frac{C_{1}}{R^{\frac{\gamma}{2}}}\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|\leq K\}}\frac{|u_{k}|^{\tau}}{|x|^{\frac{\gamma}{2}}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\frac{C_{2}}{R^{\frac{\gamma}{2}}}\sup_{k\in\mathbb{N}}\|u_{k}\|^{\tau}+\int_{B_{R}^{c}}\frac{G(x,u)}{|x|^{\gamma}}\,\mathrm{d}x,$
where $C_{2}>0$ is a suitable constant independent on $k$. The above inequality with (1.2), Corollary 3.1, Theorem 1.7 and the Lebesgue dominated convergence theorem implies
$\lim_{K\to\infty}\lim_{R\to\infty}\lim_{k\to\infty}\int_{B_{R}^{c}\cap\{x\in\mathbb{R}^{n}\colon~{}|u_{k}|\leq K\}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$
In conclusion, we have
$\lim_{R\to\infty}\lim_{k\to\infty}\int_{B^{c}_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.24)
In virtue of (6.23) and (6.24), we deduce that (6.21) holds. Hence we obtain from (6.17), (6.20), Lemma 6.6 and the Lebesgue dominated convergence theorem that
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta\psi\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta\psi~{}\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}\psi\,\mathrm{d}x,~{}\forall~{}\psi\in C^{\infty}_{0}(\mathbb{R}^{n}).$
On exploiting the density of $C^{\infty}_{0}(\mathbb{R}^{n})$ in $E$, it follows that $u$ is a weak solution of ($\boldsymbol{S}$).
To complete the proof, it is sufficient to show the compactness of the $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ in $E$. Invoking the fact that $c>0$, we have two possibilities as discussed below.
Case-1:$(c\neq 0,~{}u=0)$ In this situation, from (6.21), one has
$\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\to 0\quad\text{as}\quad k\to\infty.$
This together with $J(u_{k})\to c$ as $k\to\infty$ yields
$c+o_{k}(1)$ $=\frac{1}{p}\|\Delta u_{k}\|^{p}_{p}+\frac{2}{n}\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\geq\frac{2}{n}\big{(}\|\Delta u_{k}\|^{p}_{p}+\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\big{)}$
$\geq\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\Big{(}\|\Delta u_{k}\|_{p}^{p}+\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\Big{)}\quad\text{as}\quad k\to\infty.$
The above inequality together with $c\in(0,c_{0})$ implies the validity of (6.4), we thank to Lemma 6.5. Hence, up to a subsequence, still not relabelled, we can assume that
$\sup_{k\in\mathbb{N}}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}.$
Let us fix $m\in(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}})$ and choose $r$, $r^{\prime}$ such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Pick $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that $\alpha rm<\beta_{\gamma,n}$. Therefore, by using (1.1), Theorem 1.3, Corollary 2.6, Theorem 1.7 and Hölder’s inequality, we have
$\bigg{|}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{|}$ $\leq\zeta\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x+D_{\zeta}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha r,j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\lesssim\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\beta_{\gamma,n},j_{0}}(u_{k}/\|u_{k}\|)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\lesssim\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}\to 0\quad\text{as}\quad k\to\infty.$
It follows that
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.25)
Hence we deduce from $\langle{J^{\prime}(u_{k}),u_{k}}\rangle\to 0$ as $k\to\infty$ and (6.25) that $u_{k}\to 0$ in $E$ as $k\to\infty$. Consequently, since $J\in C^{1}(E,\mathbb{R})$, one can easily obtain $c=0$, which is a contradiction. Thus this situation is impossible.
Case-2:$(c\neq 0,~{}u\neq 0)$ In view of $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$, we obtain from the lower semicontinuity of the norm that, up to a subsequence, still denoted by iteslf, such that $\|u\|\leq\lim_{k\to\infty}\|u_{k}\|$. Indeed, if $\|u\|=\lim_{k\to\infty}\|u_{k}\|$, then by using the result of Autuori-Pucci [[6], Corollary A.2], one has $u_{k}\to u$ in $E$ as $k\to\infty$. It follows that the sequence $\{u_{k}\}_{k}$ satisfies the $(PS)_{c}$ compactness condition and thus $u$ is a nontrivial mountain pass solution of ($\boldsymbol{S}$), thanks to $J\in C^{1}(E,\mathbb{R})$.
However, to complete the proof, we simply need to show $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ is not possible. Assume by contradiction that $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ holds true. Define the sequence $\{\psi_{k}\}_{k}$ and $\psi_{0}$ as follows
$\psi_{k}:=\frac{u_{k}}{\|u_{k}\|}\quad\text{and}\quad\psi_{0}:=\frac{u}{\lim_{k\to\infty}\|u_{k}\|}.$
In view of (6.10), it is standard to check that $\|\psi_{k}\|=1$, $\psi_{0}\neq 0$ and $\psi_{k}\rightharpoonup\psi_{0}$ in $E$ as $k\to\infty$. Moreover, if $\|\psi_{0}\|=1$, then again by using [6, Corollary A.2], we have $u_{k}\to u$ in $E$ as $k\to\infty$. This contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and the proof is finished.
Next, we consider the case $\|\psi_{0}\|<1$. By direct calculations and arguing similarly as in the proof of Lemma 6.5, up to a subsequence still denoted by itself, one sees that
$\lim_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}\big{(}1-\|\psi_{0}\|^{\frac{n}{2}}\big{)}^{-\frac{2}{n-2}}.$
Let $r$, $r^{\prime}>1$ be such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. From the above inequality, we infer that there esists $m>0$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that
$\alpha r\|u_{k}\|^{\frac{n}{n-2}}<m<\beta_{\gamma,n}\big{(}1-\|\psi_{0}\|^{\frac{n}{2}}\big{)}^{-\frac{2}{n-2}}\quad\text{for large}\quad k.$
Hence, from (1.1), Hölder’s inequality, Theorem 1.4, Theorem 1.7, Corollary 2.6, we have for large $k$ that
$\bigg{|}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})(u_{k}-u)}{|x|^{\gamma}}\,\mathrm{d}x\bigg{|}$ $\leq\zeta\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{\tau-1}{\tau}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\tau}}$
$\quad+D_{\zeta}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{(q-1)r^{\prime}}|u_{k}-u|^{r^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha r,j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\lesssim\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{\tau-1}{\tau}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\tau}}+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{q-1}{qr^{\prime}}}$
×$\quad\times\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{qr^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{m,j_{0}}(\psi_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
$\lesssim\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{\tau}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{\tau}}+\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{qr^{\prime}}}=o_{k}(1)\quad\text{as}\quad k\to\infty.$
It follows that
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})(u_{k}-u)}{|x|^{\gamma}}\,\mathrm{d}x=0.$
The above convergence together with $\langle{J^{\prime}(u_{k}),u_{k}-u}\rangle=o_{k}(1)$ as $k\to\infty$ yields
$\int_{\mathbb{R}^{n}}|\Delta u_{k}|^{p-2}\Delta u_{k}\Delta(u_{k}-u)\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u_{k}|^{\frac{n}{2}-2}\Delta u_{k}\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty.$
In addition, we obtain from $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$ that
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta(u_{k}-u)\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty.$
Due to convexity of the map $s\mapsto\frac{1}{a}|s|^{a}$ for all $a\geq 1$, the above convergences imply that
$\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty$ (6.26)
for $t\in\{p,\frac{n}{2}\}$. Recall the well known Simon’s inequality as follows
$|x-y|^{\wp}\leq\begin{cases}C_{\wp}\big{[}\big{(}|x|^{\wp-2}x-|y|^{\wp-2}y\big{)}(x-y)\big{]}^{\frac{\wp}{2}}\big{[}|x|^{\wp}+|y|^{\wp}\big{]}^{\frac{2-\wp}{2}},\quad\text{if}~{}1<\wp<2,\\ c_{\wp}\big{(}|x|^{\wp-2}x-|y|^{\wp-2}y\big{)}(x-y),\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}\wp\geq 2,\end{cases}$ (6.27)
for all $x,y\in\mathbb{R}^{d}$, where $d\geq 1$, $C_{\wp}$ and $c_{\wp}$ are positive constants depending only on Weierstrass-$\wp$. By using (6.26) and (6.27), we first obtain for any $t\in\{p,\frac{n}{2}\}\geq 2$ that
$\int_{\mathbb{R}^{n}}|\Delta u_{k}-\Delta u|^{t}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty.$
Now assume that $t=p\in(1,2)$. Then according to (6.26), (6.27) and Hölder’s inequality, one has
$\int_{\mathbb{R}^{n}}|\Delta u_{k}-\Delta u|^{t}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\Big{(}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\Big{)}^{\frac{t}{2}}\Big{(}|\Delta u_{k}|^{t}+|\Delta u|^{t}\Big{)}^{\frac{2-t}{2}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\Big{(}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\Big{)}^{\frac{t}{2}}\Big{(}\big{(}|\Delta u_{k}|^{t}\big{)}^{{}^{\frac{2-t}{2}}}+\big{(}|\Delta u|^{t}\big{)}^{{}^{\frac{2-t}{2}}}\Big{)}\,\mathrm{d}x$
$\lesssim\Bigg{(}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x\Bigg{)}^{\frac{t}{2}}\Big{(}\|\Delta u_{k}\|_{t}^{\frac{(2-t)t}{2}}+\|\Delta u\|_{t}^{\frac{(2-t)t}{2}}\Big{)}$
$\lesssim\Bigg{(}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x\Bigg{)}^{\frac{t}{2}}=o_{k}(1)\quad\text{as}\quad k\to\infty.$
In conclusion, we deduce that $\Delta u_{k}\to\Delta u$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$ and thus $u_{k}\to u$ in $E$ as $k\to\infty$. Consequently, it contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and hence the result follows. ∎
Proof of Theorem 1.8..
Let $\lambda_{0}$ be defined as in Lemma 6.4 and the hypothesis $(g_{6})$ holds for all $\lambda\geq\lambda_{0}$. Then by employing Proposition 6.7 and Theorem 6.2, the functional $J$ has a nontrivial critical point $u\in E$, which is a weak solution to ($\boldsymbol{S}$). This completes the proof. ∎
Acknowledgements
Deepak Kumar Mahanta expresses his deep appreciation for the DST INSPIRE Fellowship, reference number DST/INSPIRE/03/2019/000265, supported by the Government of India. Tuhina Mukherjee expresses gratitude for the assistance received from the CSIR-HRDG grant; sanction no. 25/0324/23/EMR-II. Abhishek Sarkar received funding from the DST-INSPIRE Grant DST/INSPIRE/04/2018/002208, funded by the Government of India.